### v1

# MIGE: A Unified Framework for Multimodal Instruction-Based Image Generation and Editing

Abstract
Despite significant progress in diffusion-based image generation, subject-driven generation and instruction-based editing remain challenging. Existing methods typically treat them separately, struggling with limited high-quality data and poor generalization. However, both tasks require capturing complex visual variations while maintaining consistency between inputs and outputs. Therefore, we propose MIGE, a unified framework that standardizes task representations using multimodal instructions. It treats subject-driven generation as creation on a blank canvas and instruction-based editing as modification of an existing image, establishing a shared input-output formulation. MIGE introduces a novel multimodal encoder that maps free-form multimodal instructions into a unified vision-language space, integrating visual and semantic features through a feature fusion mechanism. This unification enables joint training of both tasks, providing two key advantages: (1) Cross-Task Enhancement: By leveraging shared visual and semantic representations, joint training improves instruction adherence and visual consistency in both subject-driven generation and instruction-based editing. (2) Generalization: Learning in a unified format facilitates cross-task knowledge transfer, enabling MIGE to generalize to novel compositional tasks, including instruction-based subject-driven editing. Experiments show that MIGE excels in both subject-driven generation and instruction-based editing while setting a state-of-the-art in the new task of instruction-based subject-driven editing. Code and model have been publicly available at this link.
MIGE: A Unified Framework for Multimodal Instruction-Based
Image Generation and Editing
Xueyun Tian♠♡, Wei Li, Bingbing Xu, Yige Yuan♠♡, Yuanzhuo Wang, Huawei Shen♠♡ CAS Key Laboratory of AI Safety, Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China University of Chinese Academy of Sciences, Beijing, China {tianxueyun23z, xubingbing, yuanyige20z, wangyuanzhuo, shenhuawei}@ict.ac.cn    weili.ucas.ict@gmail.com
Figure 1: Demonstrating the comprehensive capabilities of MIGE. As a unified framework, MIGE excels in subject-driven generation and instruction-based editing while performing well in the new compositional task of instruction-based subject-driven editing.
1 Introduction
Recent advances in diffusion models Rombach et al. (2022); Peebles and Xie (2023); Labs (2023); Esser et al. (2024) have greatly advanced customized image generation. Central to this field are subject-driven image generation, which focuses on preserving a given subject, and instruction-based image editing, which emphasizes flexible modifications based on textual instructions.
Existing methods address these tasks separately, each tailored to a specific input-output format, limiting generalization and hindering a unified instruction-following model. Subject-driven generation aims to preserve subject identity while adhering to textual instructions. Fine-tuning methods, like DreamBooth Ruiz et al. (2023), require separate training for each subject during testing, which limits generalization. In contrast, zero-shot methods like KOSMOS-G Pan et al. (2023) rely on multimodal alignment but struggle with instruction following, mainly due to the scarcity of high-quality multimodal training data, especially for multi-subject inputs. Instruction-based editing, exemplified by InstructPix2Pix Brooks et al. (2023) and EmuEdit Sheynin et al. (2024), modifies existing images based on instructions. These methods require large, high-quality editing datasets to handle diverse instructions but struggle with maintaining local-to-global consistency.
Both tasks face significant challenges due to limited data, which hinders their ability to adapt to diverse subjects or editing requirements. Furthermore, they are typically trained independently, lacking cross-task consistency, which prevents a unified understanding of the tasks. Despite these challenges, the two tasks share a fundamental principle: preserving visual consistency between inputs and outputs while capturing visual variations based on instructions. Additionally, they also follow a similar input-output paradigm, with complementary data focuses. Subject-driven image generation trains models to produce subject-accurate images, while instruction-based editing focuses on manipulating images without altering irrelevant identities. The above limitations as well as the complementary nature underscore the need for a unified framework that combines the strengths of both tasks to achieve improved performance.
Inspired by the above analysis, we propose a unified framework, MIGE, which leverages Multimodal Instructions to unify subject-driven Generation and instruction-based Editing. This vision-language representation enables substituting both entities and entire images in text prompts with their visual counterparts, allowing flexible task and instruction combinations. Structurally, we model subject-driven generation as creating an image on a blank canvas and instruction-based editing as modifying an existing image. This coherent input-output mapping simplifies the process while enabling both tasks to reinforce each other through joint training. Moreover, this integrated approach fosters emergent compositional capabilities beyond what either task can achieve alone.
Building on our unified framework, we address two critical challenges. First, for multimodal instruction encoding, existing methods Li et al. (2023a); Pan et al. (2023); Li et al. (2024b) primarily extract semantic features from images by CLIP vision encoder Radford et al. (2021), which is insufficient for preserving fine-grained subject details. To overcome this limitation, we introduce a multimodal encoder equipped with a feature fusion mechanism. This innovation integrates VAE Kingma (2013) visual features into semantic tokens, effectively capturing both detailed visual information and semantics. Secondly, we enhance compositionality through joint training, facilitating instruction-based subject-driven editing. To further optimize performance in this complex task, we develop a novel data construction pipeline based on a Multimodal Large Language Model (MLLM). This pipeline autonomously generates diverse multimodal instructions and corresponding output images. Moreover, to address the absence of an evaluation benchmark for this new composition task, we introduce MIGEBench. This specialized benchmark evaluates compositionality in terms of subject preservation and instruction adherence, providing a comprehensive assessment of the new task.
Our experiments demonstrate that joint training under the MIGE framework enables mutual reinforcement between subject-driven generation and instruction-based editing by leveraging complementary data, leading to significant performance gains over training on individual tasks. Specifically, we observe improved instruction adherence on DreamBench Ruiz et al. (2023) and enhanced detail preservation on EmuEdit and MagicBrush Zhang et al. (2024) test sets. Furthermore, MIGE achieves state-of-the-art results on MIGEBench, showcasing its fine-grained controllability in the emerging compositional tasks. Extensive case studies (as shown in Figure 5 and Figure 6) further illustrate the ability of MIGE to generate precise and consistent outputs, highlighting its effectiveness in diverse scenarios.
In summary, our contributions include:
  • Unified Framework: We propose MIGE, a unified framework that integrates subject-driven generation and instruction-based editing, facilitating joint training and mutual enhancement of both tasks.
  • Compositional Capability: We observed that joint training unlocks the compositional capability of instruction-based subject-driven editing. We propose a novel data construction pipeline and introduce the MIGEBench benchmark for comprehensive evaluation.
  • Strong Performance: The extensive experimental results show that MIGE achieves competitive results in both subject-driven generation and instruction-based editing while establishing a new state-of-the-art in instruction-based subject-driven image editing.
2 Related Work
In this section, we focus on discussing the most relevant studies in this section. A more comprehensive discussion, including additional related work, can be found in the Appendix D.
Universal generative model
Unifying tasks in image generation is challenging, and several approaches have been proposed. Pixwizard Lin et al. (2024) leverages task vectors to integrate image generation and editing but struggles with vision-language instructions due to architectural limitations. Instruct-Imagen Hu et al. (2024) unifies tasks through multimodal instructions but remains ineffective for editing even after fine-tuning. ACE Han et al. (2024) encodes various inputs as condition units within a transformer but is limited to editing. OmniGen Xiao et al. (2024b) concatenates conditions before the noise and uses autoregressive transformers, but fails to distinguish between background and reference images, resulting in poor background preservation. UniReal Chen et al. (2024) makes finer distinctions but requires complex prompts for task differentiation. In contrast, MIGE ensures input-output consistency with conditional input and multimodal instructions, simplifying task distinction and offering better scalability without the need for complex hierarchical prompts.
Subject-Driven Image Editing
Subject-driven image editing allows image modifications based on a user-specified subject, typically through addition or replacement. Methods like DreamEdit Li et al. (2023b) and DreamMix Yang et al. (2024) require multiple input images for fine-tuning, associating new concepts with special tokens, followed by iterative inpainting. PBE Yang et al. (2023) conditions a diffusion model on compressed reference image representations, while TIGIC Li et al. (2025) integrates reference images by positioning them in predefined backgrounds. MADD He et al. (2024) offers more diverse control by providing masks, boxes, or target coordinates. Overall, these methods rely on multiple reference images or precise cues like bounding boxes and masks, and cannot follow raw textual instructions for accurate and controllable generation, making them inflexible.
Figure 2: Demonstration of MIGE as a unified framework for processing multimodal instructions and conditional inputs across diverse tasks and scenarios.
Figure 3: Overall framework of MIGE. MIGE consists of two components: a multimodal encoder for processing multimodal instructions and a transformer-based diffusion model for modeling input-output relationships. The encoder incorporates a feature fusion mechanism to integrate visual and semantic features from reference image.
3 Unified Framework
3.1 Task Unification
Current methods treat subject-driven generation and instruction-based image editing separately, hindering performance due to data limitations and poor generalization. Since both aim to preserve visual consistency while incorporating instructed changes, unifying them allows mutual improvement. Joint training on diverse data enhances subject preservation and instruction following beyond the capabilities of individual task-specific models.
Therefore, we propose MIGE to unify the two tasks for joint training. By using multimodal instructions as a unified task representation, it supports flexible combinations and provides multimodal guidance. Additionally, we use conditional input to structurally unify the tasks, enhancing visual consistency. Their combination not only provides rich visual and instructional information but also naturally represents different tasks. We demonstrate the input-output format of MIGE in Figure 2.
Unified Multimodal Instruction
To enable joint training across multiple tasks, a unified task representation is essential. We introduce multimodal instructions composed of interleaved images and text, which provide both visual references and textual guidance for various controllable generation tasks. As shown in Figure 2, “<imagehere>” serves as a placeholder, sequentially replaced by input images, which can be the reference subject or the entire scene, complementing the semantics to form interleaved expressions. This unified approach effectively accommodates subject-driven generation and instruction-based editing, while remaining extensible to the more complex combination task.
Unified Conditional Input
We adopt a conditional input design to unify tasks structurally, ensuring clear task distinction while enabling shared capabilities. By concatenating different conditions, we differentiate initial generation states and capture task-specific nuances, enhancing execution accuracy. As shown in Figure 2, for instruction-based editing, we concatenate the VAE-encoded source image with the noise tensor, guiding generation based on the given image. For subject-driven generation, we use an all-zero tensor akin to a blank canvas, prompting the model to generate freely while preserving specified visual features. This design effectively captures the differences between the two tasks while ensuring input-output consistency in editing. Moreover, structuring tasks within a unified framework allows the model to leverage shared capabilities and easily extend to new tasks.
3.2 Architecture
((a))
((b))
Figure 4: Data construction pipelines for instruction-based subject-driven image editing.
The architecture of MIGE, as shown in Figure 3, consists of two main components: a multimodal encoder for processing multimodal instructions and a transformer-based diffusion model Peebles and Xie (2023) for modeling input-output relationships. The diffusion model takes concatenated latent noise and conditional input along the channel dimension as input, performing controllable generation under the control of multimodal conditions. To further enhance the integration of visual and semantic information from the reference image, we introduce a novel feature fusion mechanism in the encoder.
Multimodal Encoder
To map the multimodal instructions into a unified vision-language semantic space, we design a multimodal encoder comprising a Large Language Model (LLM) and an image feature encoding component, which includes a pretrained VAE encoder Kingma (2013) for visual feature extraction, a pretrained ViT from EVA-CLIP Fang et al. (2023) for semantic feature extraction, a Q-Former Li et al. (2023a), and a linear projection layer. Each image is represented by 32 tokens, which are then encoded by the LLM alongside text tokens to form a unified multimodal condition.
Unlike prior works Li et al. (2024a); Pan et al. (2023); Li et al. (2024b) that mainly extract semantic features of the reference images, which lack the fine-grained details necessary for preserving subject-specific features. To address this, we propose a feature fusion mechanism that combines the strengths of different visual encoders. ViT serves as a semantic feature extractor, while a VAE encoder acts as a visual feature extractor, leveraging its ability to compress and reconstruct images. As shown in Figure 3, we use the ViT semantic features compressed by the Q-Former, denoted as $f_{s}$, as guidance to adaptively incorporate the visual features $f_{v}$ extracted by the VAE. The fusion mechanism is expressed as: $f_{\text{img}}=f_{s}+\text{MLP}(\text{Attn}(f_{s},f_{v}))$ where $\text{Attn}(f_{s},f_{v})=\frac{Q(f_{s})\cdot K(f_{v})}{\sqrt{d}}V(f_{v})$. Here $d$ denotes the output dimension of the features, $Q$ is a linear layer, and both $K$ and $V$ are two-layer networks, where $K(\cdot)=V(\cdot)=\text{Linear}(\text{GELU}(\text{Linear}(\cdot)))$.
Through this fusion mechanism, we can effectively capture both visual and semantic information of reference images simultaneously without introducing additional image tokens.
3.3 Joint Training
Multimodal instructions and conditional input unify task representation and input-output formats, enabling joint training. We fine-tune MIGE on data from all tasks to enhance cross-task synergy. Except for the two image encoders, all parameters are jointly trained to align the conditional space of the diffusion model with the multimodal encoder, as shown in Figure 3. This approach improves task coordination and consistency across modalities.
Data Construction
Joint training enables multi-task learning, balancing subject preservation and instruction, followed by modeling task relationships. We create a multi-task dataset for joint multimodal instruction tuning, covering subject-driven image generation, instruction-based image editing, and instruction-based subject-driven image generation.
For subject-driven image generation, we follow the data construction methods of KOSMOS-G Pan et al. (2023) and UNIMO-G Li et al. (2024b), using an LLM to extract entities from captions, which are then fed into Grounded SAM Ren et al. (2024) for segmentation. Subjects200k dataset from OmniControl Tan et al. (2024) is also incorporated for better object preservation. For instruction-based editing, we filter existing datasets and use a rule-based strategy to construct multimodal instructions.
Instruction-based subject-driven image generation is an emerging task that involves two subtasks: instruction-based subject addition and subject replacement. This allows users to add or replace a specified subject in an image using multimodal instructions, as shown in Figure 1. However, there is no sufficient dataset available for this task.
For instruction-based subject addition, we propose a pipeline inspired by SAM-FB He et al. (2024), as shown in Figure 4(a). Starting with the SA-1B Kirillov et al. (2023) dataset, we construct the source image and a multimodal instruction for input-output pairs. We use SAM Kirillov et al. (2023) to segment annotated entities, filter and retain the main subject using MLLM, inpaint the remaining parts to create the background, and then combine the subject name with the target image to generate the multimodal instruction via GPT-4o Hurst et al. (2024). Due to resource limitations, we only process a part of the SA-1B dataset and obtain about 200k samples, but the pipeline can be scaled to generate more. For instruction-based subject replacement, we filter from existing editing data, using Grounded SAM to obtain subject segmentation and construct the multimodal instruction to form input-output pairs, as shown in Figure 4(b). We also introduce virtual try-on data constructed using IDM-VTON Choi et al. (2024), resulting in approximately 110k data samples. More details about training data construction are in Appendix B.
Figure 5: Qualitative comparison for subject-driven image generation (top rows) and instruction-based image editing (bottom rows). We compare the universal model and task-specific models on the two tasks, respectively. The prompts listed in the figure are used for MIGE and vary according to the usage of each model.
4 Experiments
4.1 Implementation Details
MIGE includes a conditional diffusion model and a multimodal encoder. Our design allows flexible selection of various diffusion models, and we initialize with PIXART-$\alpha$ Chen et al. (2023) pretrained at a 512×512 resolution. Parameters introduced for handling conditional inputs are initialized to zero, while the original weights of the model remain unchanged. The multimodal encoder consists of a pretrained Flan-T5-XXL Chung et al. (2024) as LLM for initialization and an image encoding component. This includes query tokens, Q-Former, and a projector, all initialized with a BLIP-2 Li et al. (2023a) checkpoint (pretrain_flant5xxl). The frozen VAE encoder, used as the visual feature extractor, is the same as the one in the diffusion model. Additionally, a zero-initialized MLP layer is introduced in the feature fusing mechanism for progressive visual feature integration.
MIGE is trained on our multi-task dataset using the AdamW optimizer Loshchilov (2017) with a weight decay of 0.03 and a learning rate of 1e-5 for 18 epochs on 48 H20 GPUs, totaling six days of training with a batch size of 960 (20 per GPU). A 1:1 sampling strategy is applied for subject addition and replacement tasks, and during training, there is a 5% chance of dropping either the conditional input or multimodal condition, with an additional 5% chance of dropping both, enabling classifier-free guidance during inference.
4.2 Evaluation Results
As a unified model, MIGE excels in various image generation and editing tasks, outperforming existing task-specific models. In this section, we highlight its strong performance in subject-driven image generation and instruction-based editing, showcasing its emerging capability in instruction-based subject-driven image generation on our new benchmark. For detailed evaluation results, refer to Appendix C. Additional results on these tasks can be found in Figure 12.
4.2.1 Subject-driven Image Generation
Generating images that satisfy both image and text constraints from a multimodal prompt is a challenging task. We compare MIGE with two task-specific methods that also use MLLM to encode multimodal conditions, as well as two universal models, as shown in Table 1. DINO and CLIP-I were used to assess subject fidelity, while CLIP-T evaluated adherence to multimodal instructions. Our results on DreamBench demonstrate that MIGE achieves competitive text fidelity while better preserving subject features, particularly outperforming the DINO metric. The qualitative comparison in Figure 5 further demonstrates that MIGE not only performs better in single-subject generation tasks but also preserves the distinct characteristics of each subject in multi-subject generation, whereas other models either fail to retain all subjects or lose their individual features. This advantage arises from the ability of MIGE to flexibly combine multiple reference entities in multimodal instructions and integrate additional visual features through its feature fusion mechanism.
Methods DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
Task-specific Models
KOSMOS-G Pan et al. (2023) 0.694 0.847 0.287
UNIMO-G Li et al. (2024b) 0.668 0.841 0.329
Universal Models
Omnigen (our test) Xiao et al. (2024b) 0.711 0.800 0.312
Unireal Chen et al. (2024) 0.702 0.806 0.326
MIGE (ours) 0.744 0.830 0.293
          /only_subject data 0.726 0.823 0.289
          /wo_VAE feature 0.741 0.828 0.287
Table 1: Quantitative results for subject-driven image generation on DreamBench. MIGE outperforms universal models in subject preservation and remains competitive with task-specific models.
4.2.2 Instruction-based Image Editing
Instruction-based image editing enables users to modify a source image based on free-form multimodal instructions, including adding, removing, modifying object properties, or altering the overall style. Table 2 presents a quantitative analysis of the Emu Edit and MagicBrush test sets. DINO and CLIP-I evaluate similarity to the source image, while CLIP-T measures alignment with the target description. CLIPdir quantifies the alignment between CLIP text and image embedding changes, whereas L1 and L2 capture pixel-level differences.
As shown in Table 2, MIGE achieves the highest CLIP-T score and surpasses all task-specific models in CLIPdir, demonstrating its superior ability to effectively follow multimodal instructions. As a universal model, MIGE outperforms all other universal models across all metrics on the MagicBrush test set, achieving the lowest L1 and L2 scores and the highest CLIP-I, DINO, and CLIP-T scores, highlighting its strong instruction fidelity and fine-detail preservation. This capability is further illustrated in Figure 5, where MIGE is the only model that accurately follows the instruction to add a Daffy Duck image to the red suitcase without altering other unrelated areas.
Emu Edit Test set MagicBrush Test set
[…]
Task-specific Models
InstructPix2Pix Brooks et al. (2023) 0.759 0.831 0.288 0.086 0.122 0.036 0.763 0.843 0.289 0.105 0.097 0.028
MagicBrush Zhang et al. (2024) 0.800 0.857 0.295 0.102 0.085 0.027 0.847 0.888 0.304 0.127 0.062 0.018
UltraEdit Zhao et al. (2024) 0.862 0.875 0.301 0.100 0.050 0.008 0.879 0.897 0.305 0.119 0.042 0.006
Universal Models
PixWizard Lin et al. (2024) 0.824 0.862 0.288 0.066 0.105 0.038 0.865 0.897 0.289 0.054 0.073 0.023
ACE Han et al. (2024) 0.847 0.877 0.299 0.104 0.079 0.028 0.856 0.899 0.304 0.127 0.065 0.023
OmniGen Xiao et al. (2024b) 0.766 0.819 0.299 0.123 0.160 0.065 0.821 0.863 0.289 0.087 0.116 0.045
MIGE (ours) 0.832 0.865 0.306 0.114 0.088 0.027 0.889 0.905 0.306 0.119 0.055 0.013
         /only_edit data 0.785 0.841 0.302 0.104 0.117 0.046 0.796 0.862 0.300 0.111 0.094 0.036
         /wo_VAE feature 0.799 0.846 0.300 0.098 0.103 0.035 0.811 0.868 0.299 0.105 0.098 0.036
         /wo_multimodal instruction 0.592 0.723 0.280 0.113 0.177 0.074 0.548 0.714 0.277 0.135 0.170 0.066
Table 2: Quantitative results for instruction-based image editing on Emu Edit test set and MagicBrush test set. MIGE achieves the best overall performance on MagicBrush test set among universal models, with strong instruction fidelity and detail preservation.
Figure 6: Qualitative results on the benchmark for the subject addition and subject replacement. The upper section compares subject addition results, while the lower section compares subject replacement. During testing, the <imagehere> placeholder in the multimodal instruction is replaced according to the image sequence. MIGE demonstrates flexibility in editing and excels in subject preservation ability and input-output consistency.
4.2.3 Instruction-based Subject-driven
   Image Editing
Benchmark Construction
Instruction-based subject-driven image editing is a novel task. Existing methods rely on masks or positional coordinates for editing Li et al. (2023b); Yang et al. (2024) but lack support for instruction-based editing. Current benchmarks for subject addition and replacement separately evaluate foreground and background similarities without providing a complete edited image as ground truth, making them unsuitable for this task.
To address these issues, we construct a benchmark of 1,000 test samples: 500 for instruction-based subject addition and 500 for subject replacement. Data is sourced from SEED-Data-Edit Ge et al. (2024), with subjects extracted using Grounded SAM Ren et al. (2024). Captions for target images are generated by GPT-4o and refined through manual review. For compatibility with prior methods, our benchmark also includes masks. Details of the construction process are provided in Appendix A.
Evaluation Results
Our evaluation focuses on editing ability and subject preservation. Editing ability is assessed using DINO, CLIP-I, and CLIP-T: DINO and CLIP-I measure similarity to the ground truth image, while CLIP-T evaluates alignment with the target caption. Subject preservation is evaluated by extracting the edited subject via Grounded SAM and comparing it to the input subject image using DINO and CLIP-I. This separates the evaluation of image-level editing from subject-level feature preservation. Methods that do not support instruction-based editing are marked with a cross in the table, and masks used during testing.
Quantitative comparisons with other methods are shown in Tables 3(a) and 3(b). We computed DINO and CLIP-I metrics between the source and target images, as well as the CLIP-T metric between the source image and target caption, shown in the first row of the table (denoted as “source-target”) as a baseline. In terms of editing ability, MIGE achieves the highest overall improvement across all metrics, demonstrating its effectiveness in performing edits guided by multimodal instructions. As shown in the qualitative comparison in Figure 6, MIGE understands the meaning of “replace” in instructions, rather than simply pasting the entity onto the image. For subject preservation, the results show that MIGE achieves the best performance in both tasks, as demonstrated in Figure 6.
Methods Editing Subject Preserving
Instruction DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
source-target 0.668 0.842 0.271
PBE Yang et al. (2023) 0.810 0.885 0.304 0.521 0.792
TIGIC Li et al. (2025) 0.789 0.874 0.313 0.453 0.744
MADD He et al. (2024) 0.736 0.852 0.284 0.446 0.742
OmniGen Xiao et al. (2024b) 0.802 0.868 0.296 0.580 0.792
MIGE(ours) 0.863 0.909 0.307 0.652 0.834
     /subject data + edit data 0.789 0.873 0.299 0.503 0.764
     /only_compositional data 0.780 0.860 0.303 0.585 0.808
((a)) Results on instruction-based subject replacement.
Methods Editing Subject Preserving
Instruction DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
source-target 0.783 0.903 0.295
PBE Yang et al. (2023) 0.843 0.908 0.321 0.495 0.794
TIGIC Li et al. (2025) 0.840 0.901 0.325 0.455 0.753
MADD He et al. (2024) 0.885 0.930 0.316 0.519 0.785
OmniGen Xiao et al. (2024b) 0.791 0.870 0.312 0.605 0.814
MIGE(ours) 0.909 0.940 0.322 0.638 0.838
     /subject data + edit data 0.807 0.895 0.304 0.309 0.683
     /only_compositional data 0.879 0.928 0.324 0.577 0.820
((b)) Results on instruction-based subject addition.
Table 3: Quantitative results on instruction-based subject-driven editing. Methods marked with a cross in the instruction column use masks, while the others generate images using multimodal instructions. Overall, MIGE significantly outperforms others in both tasks, showcasing superior editing and subject preservation abilities.
4.3 Further Analysis
4.3.1 Effectiveness of Joint Training
To assess the effectiveness of joint training, we train models separately on individual datasets (denoted as “only_subject data,” “only_edit data,” and “only_compositional data”) and compare their performance with the jointly trained model. The results in Table 1 and Table 2 show that joint training leads to consistent improvements across all metrics, demonstrating that subject-driven generation and instruction-based editing reinforce each other. As shown in Table 3, joint training also enhances the performance of the compositional new task, further highlighting its overall benefits. These findings emphasize both the effectiveness and necessity of joint training. In conclusion, joint training of subject-driven generation and instruction-based editing within our unified framework not only boosts compositional capability but also improves the performance of each individual task.
4.3.2 Effectiveness of Feature Fusing
MIGE employs a feature fusion mechanism in the multimodal encoder to integrate semantic features from ViT and visual features from VAE. As shown in Table 1 and Table 2, compared to the model without VAE features (denoted as “wo_VAE feature”), incorporating VAE features significantly improves detail preservation in reference images, benefiting both subject-driven image generation and instruction-based image editing. This is particularly evident in the improved CLIP-I and DINO scores and the significant reduction in L1 and L2 metrics, demonstrating that the inclusion of additional visual features helps maintain consistency between the input and output.
4.3.3 Effectiveness of Instruction-based Subject-driven Image Editing Dataset
Joint training in subject-driven image generation and instruction-based image editing enables generalization to instruction-based subject-driven image editing (denoted as “subject data + edit data”). To enhance the capability of MIGE in this new task, particularly in understanding spatial terms and size descriptions, we constructed a task-specific dataset for joint training. As shown in Table 3(a) and Table 3(b), the task-specific data significantly improved the model’s overall abilities. This demonstrates the effectiveness of our constructed dataset, and the proposed data generation pipeline serves as a valuable reference for future dataset construction.
4.3.4 Effectiveness of Multimodal Instruction
Existing instruction-based editing works Brooks et al. (2023); Sheynin et al. (2024); Zhang et al. (2024) typically use text instructions as the conditional input, while we extend this to multimodal instructions. To measure the benefit of multimodal instructions, we trained the model with text-only editing instructions for comparison. As shown in Table 2, using multimodal instructions consistently improves performance over text-only instructions (denoted as “wo_multimodal instruction”). This enhances input-output consistency and instruction-following ability in multi-task training. The significant improvement in the L1 and L2 metrics indicates finer control over images and more accurate edits. While text-only instructions provide the necessary changes, the high CLIPdir score and lower values in other metrics show that multimodal instructions add visual context, enabling more precise and faithful modifications.
5 Conclusion
We present MIGE, a unified framework that combines subject-driven generation and instruction-based editing. By leveraging multimodal instructions and conditional input, MIGE enables joint training that reinforces both tasks. This approach improves task synergy and addresses data scarcity. Joint training also unlocks new capabilities like instruction-based subject-driven image editing. We introduce pipelines for this new task to construct training data and MIGEBench for evaluation. Our experiments show that joint training leads to significant improvements in subject fidelity and instruction adherence, demonstrating the effectiveness of unifying these tasks. This integration enhances controllability and offers promising directions for future multimodal image generation and editing.
6 Limitations
Our model faces common challenges in context accuracy, complex composition, and visual faithfulness, especially in multi-entity tasks. Ethical concerns, like deepfakes, also arise. Nevertheless, our framework shows promise in enabling unified controlled image generation. Instruction-based subject-driven generation remains challenging, with difficulties in handling spatial relationships and adjusting subject sizes to fit the background. Additionally, the technology carries the risk of generating misleading images, but our work is intended solely for research, emphasizing its exploratory nature rather than real-world deployment.
Appendix A Benchmark Construction
Instruction-based subject-driven image editing is a novel task lacking evaluation benchmarks. Previous works Li et al. (2023b); Yang et al. (2024) did not support instructions and lacked ground truth target images. To address this, we designed a benchmark construction pipeline with multiple manual inspections to ensure image and caption quality.
As shown in Figure 7, we first filter valid editing pairs using a Gradio Abid et al. (2019) interface, as shown in Figure 8. Then, Qwen2.5-14B-Instruct extracts entity names and constructs multimodal instructions. After manually reviewing object names, we use Grounding DINO Liu et al. (2025) for bounding boxes and SAM for segmentation. Cropped entities are saved with black and white backgrounds for diverse test scenarios. Prompts are detailed in Appendix E.
For evaluation, Qwen2-VL-7B generates captions that compare the generated images with their ground truth counterparts. These captions are then manually reviewed to identify and remove any unrecognized entities. This review process ensures that the generated results are accurate and reliable. The final benchmark consists of 500 samples for both subject addition and replacement tasks. By carefully avoiding artifacts from inpainting and incorporating multiple rounds of manual review, we guarantee the high quality of the data. Figures 9 and 10 display some example pairs from the benchmark to illustrate the evaluation process.
Figure 7: The pipeline of benchmark construction.
Figure 8: The annotation interface for benchmarks is built with Gradio.
Figure 9: Subject addition examples in our benchmark.
Figure 10: Subject replacement examples in our benchmark.
Appendix B Dataset Construction
To enable joint training of MIGE, we designed a series of pipelines for data construction and processing. Our dataset includes three tasks: subject-driven image generation, instruction-based image editing, and instruction-based subject-driven image editing. The proportions of the training data are shown in Figure 11. In this section, we will detail the pipelines for processing each data component and other related details.
Figure 11: Composition of training data.
Subject-driven Image Generation Data Construction
We construct multimodal instruction-image pairs by replacing entities in image captions with corresponding images. The dataset is sourced from BLIP3-GROUNDING-50M Xue et al. (2024) and internal data. For the former, we discard low-quality pairs with CLIP-T scores below 0.255. Entities are filtered based on annotated bounding boxes, retaining those with sizes between 0.05 and 0.8. After segmenting entities using SAM Kirillov et al. (2023), we further compute the CLIP-T score between “a photo of [entity]” and the segmented image to mitigate incomplete pairs. For internal data, we follow the KOSMOS-G and UNIMO-G processing pipelines, using MPT-7B-Instruct Team et al. (2023) to extract entities. The original image and entity text are input into Grounded SAM to obtain entity images, which are filtered by a CLIP-T score above 0.255 and a size ratio between 0.1 and 0.85. To enhance subject preservation, we incorporate 112,846 samples from subject200k, as in OminiControl Tan et al. (2024).
Instruction-based Image Editing Data Construction
Previous works on instruction-based editing relied solely on text, covering object-level additions, deletions, modifications, and global changes in background or style. We integrated data from InstructPix2Pix Brooks et al. (2023), UltraEdit Zhao et al. (2024), MagicBrush Zhang et al. (2024), SEED-Data-Edit Ge et al. (2024), GIER Shi et al. (2020), and HQ-Edit Hui et al. (2024), filtering out instances with target image aesthetic scores above 5.5. To enhance multimodal instruction alignment, we reformulated text-based instructions based on task type. Specifically, for instructions containing phrases like “in the image” or “to the image,” we replaced them with “in<imagehere>” or “to<imagehere>.” For others, we appended “of<imagehere>” for global edits or “in<imagehere>” for localized modifications.
Instruction-based Subject-driven Image Editing Data Construction
To enhance instruction-based subject-driven image editing, we propose dedicated data construction pipelines. For subject addition, we design a pipeline inspired by SAM-FB He et al. (2024) to construct background images, foreground entities, and multimodal instructions, as shown in Figure 4(a). We start with filtering SA-1B Kirillov et al. (2023) images with an aesthetic score above 5 and select entities where the foreground-to-image ratio is between 0.1 and 0.5. Qwen2-VL-7B Wang et al. (2024) is used to verify entity completeness. The selected entities serve as foregrounds, while the backgrounds are inpainted using LAMA Suvorov et al. (2022). We then extract foreground-related text using Qwen2-VL-7B and generate multimodal instructions with GPT-4o Hurst et al. (2024). Due to resource constraints, we process the first 500,000 SA-1B samples, yielding 193,247 multimodal instruction-foreground-background pairs.
For subject replacement, as shown in Figure 4(b), we aim at constructing foreground entities and multimodal instructions. We filter SEED-Data-Edit’s part1-unsplash based on predefined rules, retaining only replace-task data where the match score in the annotations exceeds 0.3 and the target image’s aesthetic score is above 5.5. Qwen2.5-14B-Instruct Team (2024) identifies whether the main entity is being replaced. The target entity is replaced with <imagehere> to construct a multimodal prompt. Using the annotated edited subject information, we apply Grounded SAM for segmentation, obtaining 79,693 multimodal instruction-replace entity image-source image-target image pairs. In addition, we construct 34,947 virtual try-on samples using IDM-VTON Choi et al. (2024).
Appendix C Evaluation Details
We evaluated the subject-driven generation capability of MIGE on DreamBench, which contains 750 prompts covering 30 subjects. For each prompt, four images were generated using seeds 0, 1, 2, and 3, resulting in a total of 3,000 images. Following KOSMOS-G, we selected one image per prompt. Then we extracted the subject using Grounded SAM and used it as input, aligning with the training process. Subject fidelity was assessed using DINO and CLIP-I, while CLIP-T measured adherence to multimodal instructions.
For instruction-based editing evaluation, the Emu Edit benchmark contains known issues, including incorrect image-caption pairs and duplicate source-target captions. Prior works have handled these inconsistencies differently, leading to incomparable results. Therefore, we reimplemented all currently available open-source methods on both test sets using a fixed random seed (seed=0) for consistency. Consequently, CLIP-I, L1, and DINO were computed on 3,589 Emu Edit test samples, while CLIPdir and CLIP-T were evaluated on 2,899 samples after removing problematic data.
Appendix D Related Work
Subject-Driven Image Generation
Recent advances in generative models have significantly improved image customization, enabling finer and more flexible control over generated content. Subject-driven image generation, which preserves the features of given images while creating new content, has garnered increasing attention. DreamBooth Ruiz et al. (2023) fine-tunes the entire U-Net using multiple images of the same subject, making it resource-intensive. IP-Adapter Ye et al. (2023) introduces a test-time tuning-free approach by decoupling image and text, training only additional attention layers, but struggles to balance prompt adherence and image fidelity. FastComposer Xiao et al. (2024a) and Blip-Diffusion Li et al. (2024a) leverage embeddings to condition images on text but face challenges in subject preservation and efficiency with generalized vision-language inputs. KOSMOS-G Pan et al. (2023) and UNIMO-G Li et al. (2024b) use Multimodal Large Language Models (MLLMs) as multimodal encoders for greater input flexibility but require large datasets to train task-specific encoders and alignment networks. However, these methods lack strong instruction-following capabilities, limiting their scalability to new tasks and applications.
Instruction-based Image Editing
Instruction-based image editing modifies images at the object level (e.g., adding, removing, or altering objects) or the image level (e.g., changing style or environment) based on user instructions. This approach offers greater flexibility than caption-based methods but demands stronger instruction-following capabilities. Inst-inpaint Yildirim et al. (2023) focuses on object removal, while EraseDraw Canberk et al. (2025) fine-tunes a model for object addition using data derived from removal tasks. InstructPix2Pix Brooks et al. (2023), trained on a large synthetic dataset, supports four types of edits, and MagicBrush Zhang et al. (2024) refines this capability by fine-tuning on real, manually annotated data. EmuEdit Sheynin et al. (2024) enhances diffusion-based editing with a task vector to differentiate tasks. However, these methods often struggle with maintaining input-output consistency.
Figure 12: Qualitative results of subject-driven image generation (top) , instruction-based image editing (middle), and instruction-based subject-driven image editing (bottom).
Appendix E Prompts
LLMs and MLLMs are integral to our training data and benchmark construction pipeline, and the prompts used are presented in this section.
Prompt for Qwen2-VL-7B to determine the completeness of subject in the image
Prompt: Determine if the subject in the image is complete. If it is complete and not an abstract object such as background, grass, sky, tree, stone, or part of another item, please return True. Otherwise, return False.
Table 4: Prompt design for subject completeness evaluation in an image. The model should identify whether the subject is complete, not abstract elements.
Prompt for GPT-4o to generate multimodal instruction
Prompt: The object is {object_name}. It is located in a bounding box with coordinates ({x}, {y}, {w}, {h}) on an image of size {width}x{height}. Describe its size, relative position, and relation to surrounding objects. Avoid describing the overall scene or unrelated elements.
Your response should start with “Add <imagehere> to the [position] of <imagehere>”. (The first <imagehere> indicates the object and the second indicates the image.) Keep the <imagehere> symbol in the first sentence in your reply.
Answer briefly in two sentences:
Table 5: Prompt design for GPT-4o to generate multimodal instruction for subject addition training data.
Prompt for Qwen2.5-14B-Instruct to determine whether an edit instruction pertains to the main subject of an image
Prompt: You are an assistant that determines whether an edit instruction pertains to the main subject of an image.
The main subject refers to humans or animals only. If the edit instruction is related to the main subject, respond with ’yes’.
If it pertains to background, large areas of vegetation, or environmental information, respond with ’no’.
Here are some examples:
1. replace the grass with sand $\rightarrow$ no
2. replace the trees with palm trees $\rightarrow$ no
3. replace the dirt road with a cobblestone path $\rightarrow$ no
4. replace the bird with a parrot $\rightarrow$ yes
Now, analyze the following instruction and respond accordingly: {instruction}
Table 6: Prompt design for determining if an edit instruction pertains to the main subject of the image. The model evaluates whether the instruction relates to humans, animals, or other environmental features.
Prompt for extracting the main subject and modifying the edit instruction
Prompt 1: Replace Examples
You are an assistant that determines the main subject of an edit instruction and outputs the extracted main subject along with a modified prompt.
Given an edit instruction like “replace X with Y”:
- Extract “Y” (the replacement subject).
- Output the extracted subject.
- Modify the instruction by replacing “Y” with <imagehere> in the original prompt and output the new instruction.
Examples:
1. replace one woman with a man $\rightarrow$ Output: a man, replace one woman with <imagehere>
2. replace the castle with another castle $\rightarrow$ Output: castle, replace the castle with <imagehere>
3. replace the desk with a white one in the bottom middle $\rightarrow$ Output: desk, replace the desk with <imagehere> in the bottom middle
Prompt 2: Add Examples
You are an assistant that extracts the added subject in the edit instruction.
Given an edit instruction like “add X to the image”:
- Extract “X” (the added subject).
- Output the extracted subject.
- Modify the instruction by replacing “X” with <imagehere> in the original prompt and output the new instruction.
Here are some examples:
1. add human over the stone in the bottom left $\rightarrow$ Output: human, add <imagehere> over the stone in the bottom left
2. add a car on the road at the bottom $\rightarrow$ Output: a car, add <imagehere> on the road at the bottom
3. Add an owl on the left shoulder $\rightarrow$ Output: an owl, Add <imagehere> on the left shoulder
Table 7: Prompt designs for extracting and modifying edit instructions based on either replacement or addition of subjects.

### v3

@@
-MIGE: A Unified Framework for Multimodal Instruction-Based Image Generation and Editing
-Abstract
+MIGE: Mutually Enhanced Multimodal Instruction-Based Image Generation and Editing
+Abstract.
 Despite significant progress in diffusion-based image generation, subject-driven generation and instruction-based editing remain challenging.
@@
 However, both tasks require capturing complex visual variations while maintaining consistency between inputs and outputs.
-Therefore, we propose MIGE, a unified framework that standardizes task representations using multimodal instructions.
-It treats subject-driven generation as creation on a blank canvas and instruction-based editing as modification of an existing image, establishing a shared input-output formulation.
-MIGE introduces a novel multimodal encoder that maps free-form multimodal instructions into a unified vision-language space, integrating visual and semantic features through a feature fusion mechanism.
+Inspired by this, we propose MIGE, a unified framework that standardizes task representations using multimodal instructions.
+It first treats subject-driven generation as creation on a blank canvas and instruction-based editing as modification of an existing image, establishing a shared input-output formulation, then introduces a novel multimodal encoder that maps free-form multimodal instructions into a unified vision-language space, integrating visual and semantic features through a feature fusion mechanism.
 This unification enables joint training of both tasks, providing two key advantages: (1) Cross-Task Enhancement: by leveraging shared visual and semantic representations, joint training improves instruction adherence and visual consistency in both subject-driven generation and instruction-based editing.
 (2) Generalization: learning in a unified format facilitates cross-task knowledge transfer, enabling MIGE to generalize to novel compositional tasks, including instruction-based subject-driven editing.
-Experiments show that MIGE excels in both subject-driven generation and instruction-based editing while setting a state-of-the-art in the new task of instruction-based subject-driven editing.
+Experiments show that MIGE excels in both subject-driven generation and instruction-based editing while setting a SOTA in the new task of instruction-based subject-driven editing.
 Code and model have been publicly available at this link.
-MIGE: A Unified Framework for Multimodal Instruction-Based
-Image Generation and Editing
-Xueyun Tian♠♡, Wei Li, Bingbing Xu, Yige Yuan♠♡, Yuanzhuo Wang, Huawei Shen♠♡ CAS Key Laboratory of AI Safety, Institute of Computing Technology, Chinese Academy of Sciences, Beijing, China University of Chinese Academy of Sciences, Beijing, China {tianxueyun23z, xubingbing, yuanyige20z, wangyuanzhuo, shenhuawei}@ict.ac.cn    weili.ucas.ict@gmail.com
-Figure 1: Demonstrating the comprehensive capabilities of MIGE.
-As a unified framework, MIGE excels in subject-driven generation and instruction-based editing while performing well in the new compositional task of instruction-based subject-driven editing.
-1 Introduction
-Recent advances in diffusion models Rombach et al.
-(2022); Peebles and Xie (2023); Labs (2023); Esser et al.
-(2024) have greatly advanced customized image generation.
-Central to this field are subject-driven image generation, which focuses on preserving a given subject, and instruction-based image editing, which emphasizes flexible modifications based on textual instructions.
-Existing methods address these tasks separately, each tailored to a specific input-output format, limiting generalization and hindering a unified instruction-following model.
-Subject-driven generation aims to preserve subject identity while adhering to textual instructions.
-Fine-tuning methods, like DreamBooth Ruiz et al.
-(2023), require separate training for each subject during testing, which limits generalization.
-In contrast, zero-shot methods like KOSMOS-G Pan et al.
-(2023) rely on multimodal alignment but struggle with instruction following, mainly due to the scarcity of high-quality multimodal training data, especially for multi-subject inputs.
-Instruction-based editing, exemplified by InstructPix2Pix Brooks et al.
-(2023) and EmuEdit Sheynin et al.
-(2024), modifies existing images based on instructions.
+Controllable Image Generation, Diffusion Model, Joint Training
+copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33th ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06submissionid: 6167ccs: Computing methodologies Image representations
+1. Introduction
+Recent advances in diffusion models(Rombach et al., 2022; Peebles and Xie, 2023; Labs, 2023; Esser et al., 2024) have significantly advanced the development of customized image generation, enabling its widespread adoption in both subject-driven generation(Ruiz et al., 2023; Ye et al., 2023; Li et al., 2024a; Pan et al., 2023; Kang et al., 2025; Kim et al., 2024; Wu et al., 2025; Liang et al., 2025; Peng et al., 2024) and instruction-based editing(Brooks et al., 2023; Sheynin et al., 2024; Fu et al., 2023; Han et al., 2024).
+The former focuses on preserving a given subject, while the latter emphasizes flexible modifications guided by textual instructions.
+Their capability to generate personalized and controllable results demonstrates substantial practical value, thereby attracting growing research attention.
+Existing approaches (Li et al., 2024a; Pan et al., 2023; Hu et al., 2024; Huang et al., 2024b) typically treat these two tasks independently and address them separately by tailoring specific input-output formats to each.
+However, this results in limited generalization capability and necessitates extensive training data.
+Specifically, subject-driven generation aims to preserve subject identity while adhering to textual instructions.
+Fine-tuning methods, like DreamBooth (Ruiz et al., 2023), require separate training for each subject during testing, which limits generalization.
+In contrast, zero-shot methods like KOSMOS-G (Pan et al., 2023) rely on multimodal alignment but struggle with instruction following, mainly due to the scarcity of high-quality multimodal training data, especially for multi-subject inputs.
+Instruction-based editing, exemplified by InstructPix2Pix (Brooks et al., 2023), modifies images based on instructions.
 These methods require large, high-quality editing datasets to handle diverse instructions but struggle with maintaining local-to-global consistency.
-Both tasks face significant challenges due to limited data, which hinders their ability to adapt to diverse subjects or editing requirements.
-Furthermore, they are typically trained independently, lacking cross-task consistency, which prevents a unified understanding of the tasks.
-Despite these challenges, the two tasks share a fundamental principle: preserving visual consistency between inputs and outputs while capturing visual variations based on instructions.
-Additionally, they also follow a similar input-output paradigm, with complementary data focuses.
-Subject-driven image generation trains models to produce subject-accurate images, while instruction-based editing focuses on manipulating images without altering irrelevant identities.
-The above limitations as well as the complementary nature underscore the need for a unified framework that combines the strengths of both tasks to achieve improved performance.
-Inspired by the above analysis, we propose a unified framework, MIGE, which leverages Multimodal Instructions to unify subject-driven Generation and instruction-based Editing.
-This vision-language representation enables substituting both entities and entire images in text prompts with their visual counterparts, allowing flexible task and instruction combinations.
+The above separate processing methods can only rely on limited data and single-task optimization objectives, which hinders their ability to adapt to diverse subjects or editing requirements, and lack cross-task consistency, preventing a unified understanding of the tasks.
+To address these challenges, we shift our perspective from separate processing to the objectives of these two tasks.
+These two tasks share a fundamental principle: maintaining visual consistency between inputs and outputs while capturing variations guided by instructions.
+They also follow a similar input-output paradigm with complementarities: subject-driven generation emphasizes subject accuracy, while instruction-based editing manipulates images without altering irrelevant identities.
+These limitations and complementarities highlight the need for a unified framework that takes advantage of the strengths of both tasks to improve performance.
+Inspired by the above analysis, we propose a unified framework, MIGE, which leverages Multimodal Instructions to unify subject-driven image Generation and instruction-based Editing.
+The vision-language representation enables substituting both entities and entire images in text prompts with their visual counterparts, allowing flexible task and instruction combinations.
 Structurally, we model subject-driven generation as creating an image on a blank canvas and instruction-based editing as modifying an existing image.
@@
 Building on our unified framework, we address two critical challenges.
-First, for multimodal instruction encoding, existing methods Li et al.
-(2023a); Pan et al.
-(2023); Li et al.
-(2024b) primarily extract semantic features from images by CLIP vision encoder Radford et al.
-(2021), which is insufficient for preserving fine-grained subject details.
-To overcome this limitation, we introduce a multimodal encoder equipped with a feature fusion mechanism.
-This innovation integrates VAE Kingma (2013) visual features into semantic tokens, effectively capturing both detailed visual information and semantics.
+First, for multimodal instruction encoding, existing methods (Li et al., 2023b; Pan et al., 2023; Li et al., 2024c) mainly rely on CLIP vision encoders (Radford et al., 2021) to extract semantic features, which is insufficient for preserving fine-grained subject details.
+To overcome this limitation, we propose a multimodal encoder with a feature fusion mechanism that integrates VAE (Kingma, 2013) visual features into semantic tokens, effectively capturing both detailed visual information and semantics.
 Secondly, we enhance compositionality through joint training, facilitating instruction-based subject-driven editing.
@@
 This specialized benchmark evaluates compositionality in terms of subject preservation and instruction adherence, providing a comprehensive assessment of the new task.
-Our experiments demonstrate that joint training under the MIGE framework enables mutual reinforcement between subject-driven generation and instruction-based editing by leveraging complementary data, leading to significant performance gains over training on individual tasks.
-Specifically, we observe improved instruction adherence on DreamBench Ruiz et al.
-(2023) and enhanced detail preservation on EmuEdit and MagicBrush Zhang et al.
-(2024) test sets.
-Furthermore, MIGE achieves state-of-the-art results on MIGEBench, showcasing its fine-grained controllability in the emerging compositional tasks.
-Extensive case studies (as shown in Figure 5 and Figure 6) further illustrate the ability of MIGE to generate precise and consistent outputs, highlighting its effectiveness in diverse scenarios.
+Our experiments demonstrate that joint training under the MIGE framework enables mutual enhancement between subject-driven generation and instruction-based editing, leading to significant performance gains over training on individual tasks.
+As shown in Figure 2, joint training markedly improves input-output consistency across both tasks, especially in editing, with scores rising from 0.821 to 0.873.
+Moreover, MIGE achieves state-of-the-art results on MIGEBench, showcasing its fine-grained controllability in emerging compositional tasks.
+Extensive case studies (Figures 78) further demonstrate the ability of MIGE to generate precise and consistent outputs, highlighting its effectiveness across diverse scenarios.
+Figure 2.
+Joint training significantly improves input-output similarity across both tasks.
+We show the average DINO and CLIP-I scores for subject-driven generation and instruction-based editing under single-task and joint training.
 In summary, our contributions include:
-  • Unified Framework: We propose MIGE, a unified framework that integrates subject-driven generation and instruction-based editing, facilitating joint training and mutual enhancement of both tasks.
-  • Compositional Capability: We observed that joint training unlocks the compositional capability of instruction-based subject-driven editing.
-We propose a novel data construction pipeline and introduce the MIGEBench benchmark for comprehensive evaluation.
-  • Strong Performance: The extensive experimental results show that MIGE achieves competitive results in both subject-driven generation and instruction-based editing while establishing a new state-of-the-art in instruction-based subject-driven image editing.
-2 Related Work
-In this section, we focus on discussing the most relevant studies in this section. A more comprehensive discussion, including additional related work, can be found in the Appendix D.
+  • Promising perspective of task mutual enhancement: We shift our perspective from traditional separate processing to mutual enhancement.
+Then, we design a unified input-output format for joint training on subject-driven generation and instruction-based editing, validating that joint training enables task enhancement.
+  • Novel framework for Compositional Capability: We propose MIGE, a unified framework integrating two mainstream tasks in controllable image generation.
+Joint training further enables the compositional capability of instruction-based subject-driven editing.
+We also introduce a novel data pipeline and MIGEBench for comprehensive evaluation.
+  • Strong Performance with Extensive Experiments: Extensive results show that MIGE performs competitively on subject-driven generation and instruction-based editing, with scores rising from 0.821 to 0.873, and sets a new state-of-the-art in instruction-based subject-driven editing.
+2. Related Work
+In this section, we discuss the most relevant studies. A more comprehensive discussion can be found in the Supplementary Material.
 Universal generative model
-Unifying tasks in image generation is challenging, and several approaches have been proposed.
-Pixwizard Lin et al.
-(2024) leverages task vectors to integrate image generation and editing but struggles with vision-language instructions due to architectural limitations.
-Instruct-Imagen Hu et al.
-(2024) unifies tasks through multimodal instructions but remains ineffective for editing even after fine-tuning.
-ACE Han et al.
-(2024) encodes various inputs as condition units within a transformer but is limited to editing.
-OmniGen Xiao et al.
-(2024b) concatenates conditions before the noise and uses autoregressive transformers, but fails to distinguish between background and reference images, resulting in poor background preservation.
-UniReal Chen et al.
-(2024) makes finer distinctions but requires complex prompts for task differentiation.
-In contrast, MIGE ensures input-output consistency with conditional input and multimodal instructions, simplifying task distinction and offering better scalability without the need for complex hierarchical prompts.
+Unifying tasks in image generation remains challenging.
+Some methods (Li et al., 2025a; Huang et al., 2024a) focus on bridging image understanding and generation, such as SEED-X (Ge et al., 2024), Emu2 (Sun et al., 2024), and WeGen (Huang et al., 2025), which use autoregressive models with auxiliary diffusion modules.
+UnifiedMLLM (Li et al., 2024b) employs MLLMs to learn shared representations and route to task-specific models.
+However, these approaches either emphasize understanding over generation or rely on complex pipelines.
+Others aim to unify generation and editing.
+Pixwizard (Lin et al., 2024) uses text-only instructions and task vectors.
+Instruct-Imagen (Hu et al., 2024) supports multimodal inputs but performs poorly on editing.
+ACE (Han et al., 2024) is limited to editing via transformer-based conditioning.
+OmniGen (Xiao et al., 2024) cannot distinguish reference images.
+OmniControl (Tan et al., 2024) merges all conditions with noise, while UniReal (Chen et al., 2024c) adds finer control but requires complex prompts.
+RealGeneral (Lin et al., 2025) enhances attention, and OnePic (Tao et al., 2024) depends on multiple references and masks.
+Overall, prior work often sacrifices flexibility or generality, combining tasks without leveraging their intrinsic connections.
+In contrast, we propose MIGE, which jointly trains on subject-driven generation and instruction-based editing, revealing their mutual benefits.
 Subject-Driven Image Editing
-Subject-driven image editing allows image modifications based on a user-specified subject, typically through addition or replacement.
-Methods like DreamEdit Li et al.
-(2023b) and DreamMix Yang et al.
-(2024) require multiple input images for fine-tuning, associating new concepts with special tokens, followed by iterative inpainting.
-PBE Yang et al.
-(2023) conditions a diffusion model on compressed reference image representations, while TIGIC Li et al.
-(2025) integrates reference images by positioning them in predefined backgrounds.
-MADD He et al.
-(2024) offers more diverse control by providing masks, boxes, or target coordinates.
-Overall, these methods rely on multiple reference images or precise cues like bounding boxes and masks, and cannot follow raw textual instructions for accurate and controllable generation, making them inflexible.
-Figure 2: Demonstration of MIGE as a unified framework for processing multimodal instructions and conditional inputs across diverse tasks and scenarios.
-Figure 3: Overall framework of MIGE.
+Subject-driven image editing enables modifications based on a user-specified subject, typically including addition or replacement.
+Existing methods often rely on multiple reference images or precise cues.
+DreamEdit (Li et al., 2023a) and DreamMix (Yang et al., 2024) require multiple images for fine-tuning and iterative inpainting.
+PBE (Yang et al., 2023) uses compressed reference embeddings, while TIGIC (Li et al., 2025c) embeds subjects into predefined backgrounds.
+MADD (He et al., 2024a) supports diverse controls via masks or coordinates.
+AnyDoor (Chen et al., 2024b), AnyEdit (Yu et al., 2024), MimicBrush (Chen et al., 2024a) and BlobCtrl (Li et al., 2025b) require masks or additional range blobs.
+FreeEdit (He et al., 2024b) is limited to instruction templates and class-aware features.
+Overall, these approaches depend on detailed inputs and lack flexibility in handling diverse instructions for controllable generation.
+Figure 3. Demonstration of MIGE as a unified framework for processing multimodal instructions and conditional inputs across diverse scenarios.
+Figure 4.
+Overall framework of MIGE.
 MIGE consists of two components: a multimodal encoder for processing multimodal instructions and a transformer-based diffusion model for modeling input-output relationships.
 The encoder incorporates a feature fusion mechanism to integrate visual and semantic features from reference image.
-3 Unified Framework
-3.1 Task Unification
+3. Method
+In this section, we first shift our view from separate training to mutual enhancement.
+Guided by this new perspective, we first introduce the unified framework MIGE, followed by details of the architecture design and joint training approach.
+3.1. Unified framework MIGE
 Current methods treat subject-driven generation and instruction-based image editing separately, hindering performance due to data limitations and poor generalization.
@@
 We demonstrate the input-output format of MIGE in Figure 3.
-Unified Multimodal Instruction
+3.1.1. Unified Multimodal Instruction
 To enable joint training across multiple tasks, a unified task representation is essential.
 We introduce multimodal instructions composed of interleaved images and text, which provide both visual references and textual guidance for various controllable generation tasks.
-As shown in Figure 2, “<imagehere>” serves as a placeholder, sequentially replaced by input images, which can be the reference subject or the entire scene, complementing the semantics to form interleaved expressions.
-This unified approach effectively accommodates subject-driven generation and instruction-based editing, while remaining extensible to the more complex combination task.
-Unified Conditional Input
-We adopt a conditional input design to unify tasks structurally, ensuring clear task distinction while enabling shared capabilities.
-By concatenating different conditions, we differentiate initial generation states and capture task-specific nuances, enhancing execution accuracy.
-As shown in Figure 2, for instruction-based editing, we concatenate the VAE-encoded source image with the noise tensor, guiding generation based on the given image.
-For subject-driven generation, we use an all-zero tensor akin to a blank canvas, prompting the model to generate freely while preserving specified visual features.
-This design effectively captures the differences between the two tasks while ensuring input-output consistency in editing.
-Moreover, structuring tasks within a unified framework allows the model to leverage shared capabilities and easily extend to new tasks.
-3.2 Architecture
-((a))
-((b))
-Figure 4: Data construction pipelines for instruction-based subject-driven image editing.
-The architecture of MIGE, as shown in Figure 3, consists of two main components: a multimodal encoder for processing multimodal instructions and a transformer-based diffusion model Peebles and Xie (2023) for modeling input-output relationships.
+As shown in Figure 3, “¡imagehere¿” serves as a placeholder, sequentially replaced by input images, which can be the reference subject or the entire scene, complementing the semantics to form interleaved expressions.
+This unified approach effectively accommodates subject-driven generation and instruction-based editing while remaining extensible to the more complex combination task.
+3.1.2. Unified Conditional Input
+We design a unified conditional input format to distinguish tasks while enabling shared capabilities.
+As shown in Figure 3, conditional inputs vary across tasks.
+For instruction-based editing, the VAE-encoded source image is used to preserve fine details of the original image, inform the model to edit only the target region, as done in prior editing methods (Brooks et al., 2023; Zhang et al., 2024).
+To standardize the input format, we concatenate an all-zero tensor for subject-driven generation to indicate generation from scratch with visual constraints from the multimodal instruction.
+This design unifies the input format, ensures input-output consistency, and facilitates extensibility to new tasks within a unified framework.
+3.2. Architecture
+The architecture of MIGE, as shown in Figure 4, consists of two main components: a multimodal encoder for processing multimodal instructions and a transformer-based diffusion model (Peebles and Xie, 2023) for modeling input-output relationships.
 The diffusion model takes concatenated latent noise and conditional input along the channel dimension as input, performing controllable generation under the control of multimodal conditions.
 To further enhance the integration of visual and semantic information from the reference image, we introduce a novel feature fusion mechanism in the encoder.
-Multimodal Encoder
-To map the multimodal instructions into a unified vision-language semantic space, we design a multimodal encoder comprising a Large Language Model (LLM) and an image feature encoding component, which includes a pretrained VAE encoder Kingma (2013) for visual feature extraction, a pretrained ViT from EVA-CLIP Fang et al.
-(2023) for semantic feature extraction, a Q-Former Li et al.
-(2023a), and a linear projection layer.
+3.2.1. Multimodal Encoder
+To map the multimodal instructions into a unified vision-language semantic space, we design a multimodal encoder comprising a Large Language Model (LLM) and an image feature encoding component, which includes a pretrained VAE encoder (Kingma, 2013) for visual feature extraction, a pretrained ViT from EVA-CLIP (Fang et al., 2023) for semantic feature extraction, a Q-Former (Li et al., 2023b), and a linear projection layer.
 Each image is represented by 32 tokens, which are then encoded by the LLM alongside text tokens to form a unified multimodal condition.
-Unlike prior works Li et al.
-(2024a); Pan et al.
-(2023); Li et al.
-(2024b) that mainly extract semantic features of the reference images, which lack the fine-grained details necessary for preserving subject-specific features.
+Prior works (Li et al., 2024a; Pan et al., 2023; Li et al., 2024c) focus on extracting semantics from reference images but lack the fine-grained details needed to preserve subject-specific features.
 To address this, we propose a feature fusion mechanism that combines the strengths of different visual encoders.
@@
 As shown in Figure 4, we use the ViT semantic features compressed by the Q-Former, denoted as $f_{s}$, as guidance to adaptively incorporate the visual features $f_{v}$ extracted by the VAE.
+Inspired by TokenPacker (Li et al., 2024d), we represent rich features with fewer tokens, and apply ReLU (Glorot et al., 2011) to enhance stability.
 The fusion mechanism is expressed as: $f_{\text{img}}=f_{s}+\text{MLP}(\text{Attn}(f_{s},f_{v}))$ where $\text{Attn}(f_{s},f_{v})=\frac{Q(f_{s})\cdot K(f_{v})}{\sqrt{d}}V(f_{v})$.
 Here $d$ denotes the output dimension of the features, $Q$ is a linear layer, and both $K$ and $V$ are two-layer networks, where $K(\cdot)=V(\cdot)=\text{Linear}(\text{GELU}(\text{Linear}(\cdot)))$.
-Through this fusion mechanism, we can effectively capture both visual and semantic information of reference images simultaneously without introducing additional image tokens.
-3.3 Joint Training
+This fusion mechanism captures both semantic and visual information without introducing extra image tokens.
+The multimodal encoder compresses each image into 32 fused tokens, which replace the “¡imagehere¿” placeholder in the interleaved instruction.
+The LLM then encodes the entire prompt into a unified vision-language representation to guide the diffusion model.
+(a) Subject Addition Dataset Construction Pipeline
+(b) Subject Replacement Dataset Construction Pipeline
+Figure 5. Data construction pipelines for instruction-based subject-driven image editing.
+3.3. Joint Training
 Multimodal instructions and conditional input unify task representation and input-output formats, enabling joint training.
@@
 This approach improves task coordination and consistency across modalities.
-Data Construction
+3.3.1. Data Construction
 Joint training enables multi-task learning, balancing subject preservation and instruction, followed by modeling task relationships.
 We create a multi-task dataset for joint multimodal instruction tuning, covering subject-driven image generation, instruction-based image editing, and instruction-based subject-driven image generation.
-For subject-driven image generation, we follow the data construction methods of KOSMOS-G Pan et al.
-(2023) and UNIMO-G Li et al.
-(2024b), using an LLM to extract entities from captions, which are then fed into Grounded SAM Ren et al.
-(2024) for segmentation.
-Subjects200k dataset from OmniControl Tan et al.
-(2024) is also incorporated for better object preservation.
+For subject-driven image generation, we follow the data construction methods of KOSMOS-G (Pan et al., 2023) and UNIMO-G (Li et al., 2024c), using an LLM to extract entities from captions, which are then fed into Grounded SAM (Ren et al., 2024) for segmentation.
+Subjects200k dataset from OmniControl (Tan et al., 2024) is also incorporated for better object preservation.
 For instruction-based editing, we filter existing datasets and use a rule-based strategy to construct multimodal instructions.
-Instruction-based subject-driven image generation is an emerging task that involves two subtasks: instruction-based subject addition and subject replacement.
-This allows users to add or replace a specified subject in an image using multimodal instructions, as shown in Figure 1.
+Instruction-based subject-driven image generation is an emerging task involving two subtasks: instruction-based subject addition and subject replacement, which allows users to add or replace a specified subject in an image using multimodal instructions.
 However, there is no sufficient dataset available for this task.
-For instruction-based subject addition, we propose a pipeline inspired by SAM-FB He et al.
-(2024), as shown in Figure 4(a).
-Starting with the SA-1B Kirillov et al.
-(2023) dataset, we construct the source image and a multimodal instruction for input-output pairs.
-We use SAM Kirillov et al.
-(2023) to segment annotated entities, filter and retain the main subject using MLLM, inpaint the remaining parts to create the background, and then combine the subject name with the target image to generate the multimodal instruction via GPT-4o Hurst et al.
-(2024).
-Due to resource limitations, we only process a part of the SA-1B dataset and obtain about 200k samples, but the pipeline can be scaled to generate more.
-For instruction-based subject replacement, we filter from existing editing data, using Grounded SAM to obtain subject segmentation and construct the multimodal instruction to form input-output pairs, as shown in Figure 4(b).
-We also introduce virtual try-on data constructed using IDM-VTON Choi et al.
-(2024), resulting in approximately 110k data samples.
-More details about training data construction are in Appendix B.
-Figure 5: Qualitative comparison for subject-driven image generation (top rows) and instruction-based image editing (bottom rows).
-We compare the universal model and task-specific models on the two tasks, respectively.
-The prompts listed in the figure are used for MIGE and vary according to the usage of each model.
-4 Experiments
-4.1 Implementation Details
+For instruction-based subject addition, we propose a pipeline inspired by SAM-FB (He et al., 2024a), as shown in Figure 5(a).
+Starting with the SA-1B (Kirillov et al., 2023) dataset, we construct the source image and a multimodal instruction for input-output pairs.
+We use SAM (Kirillov et al., 2023) to segment annotated entities, filter and retain the main subject using an MLLM, inpaint the remaining parts to create the background, and then combine the subject name with the target image to generate the multimodal instruction via GPT-4o (Hurst et al., 2024).
+Due to resource limitations, we process a part of the SA-1B dataset and obtain about 200k samples, but the pipeline can be scaled to generate more.
+For instruction-based subject replacement, we filter from existing editing data, use Grounded SAM for subject segmentation, and construct multimodal instructions to form input-output pairs, as shown in Figure 5(b).
+We also introduce virtual try-on data constructed using IDM-VTON (Choi et al., 2024), resulting in approximately 110k samples.
+More details on training data construction are in Supplementary Material.
+4. Experiments
+In this section, we present experimental results on three tasks, followed by analysis of the joint training strategy and framework design.
+For subject-driven image generation, we compare our method with both task-specific models using MLLMs as encoders and recent unified frameworks.
+For instruction-based image editing, we similarly compare against both task-specific and unified models, including SmartEdit (Huang et al., 2024b) and ml-MGIE (Fu et al., 2023), which also use MLLMs.
+For instruction-based subject-driven image editing, given the lack of instruction-only baselines, we additionally evaluate mask-dependent methods using masks from MIGEBench.
+4.1. Implementation Details
 MIGE includes a conditional diffusion model and a multimodal encoder.
-Our design allows flexible selection of various diffusion models, and we initialize with PIXART-$\alpha$ Chen et al.
-(2023) pretrained at a 512×512 resolution.
-Parameters introduced for handling conditional inputs are initialized to zero, while the original weights of the model remain unchanged.
-The multimodal encoder consists of a pretrained Flan-T5-XXL Chung et al.
-(2024) as LLM for initialization and an image encoding component.
-This includes query tokens, Q-Former, and a projector, all initialized with a BLIP-2 Li et al.
-(2023a) checkpoint (pretrain_flant5xxl).
+Our design allows flexible selection of various diffusion models, and we initialize with PIXART-$\alpha$ (Chen et al., 2023) pretrained at a 512×512 resolution.
+The multimodal encoder consists of a pretrained Flan-T5-XXL (Chung et al., 2024) as LLM for initialization and an image encoding component.
+This includes query tokens, Q-Former, and a projector, all initialized with a BLIP-2 (Li et al., 2023b) checkpoint (pretrain_flant5xxl).
 The frozen VAE encoder, used as the visual feature extractor, is the same as the one in the diffusion model.
-Additionally, a zero-initialized MLP layer is introduced in the feature fusing mechanism for progressive visual feature integration.
-MIGE is trained on our multi-task dataset using the AdamW optimizer Loshchilov (2017) with a weight decay of 0.03 and a learning rate of 1e-5 for 18 epochs on 48 H20 GPUs, totaling six days of training with a batch size of 960 (20 per GPU).
-A 1:1 sampling strategy is applied for subject addition and replacement tasks, and during training, there is a 5% chance of dropping either the conditional input or multimodal condition, with an additional 5% chance of dropping both, enabling classifier-free guidance during inference.
-4.2 Evaluation Results
-As a unified model, MIGE excels in various image generation and editing tasks, outperforming existing task-specific models.
-In this section, we highlight its strong performance in subject-driven image generation and instruction-based editing, showcasing its emerging capability in instruction-based subject-driven image generation on our new benchmark.
-For detailed evaluation results, refer to Appendix C.
-Additional results on these tasks can be found in Figure 12.
-4.2.1 Subject-driven Image Generation
-Generating images that satisfy both image and text constraints from a multimodal prompt is a challenging task.
-We compare MIGE with two task-specific methods that also use MLLM to encode multimodal conditions, as well as two universal models, as shown in Table 1.
-DINO and CLIP-I were used to assess subject fidelity, while CLIP-T evaluated adherence to multimodal instructions.
-Our results on DreamBench demonstrate that MIGE achieves competitive text fidelity while better preserving subject features, particularly outperforming the DINO metric.
-The qualitative comparison in Figure 5 further demonstrates that MIGE not only performs better in single-subject generation tasks but also preserves the distinct characteristics of each subject in multi-subject generation, whereas other models either fail to retain all subjects or lose their individual features.
-This advantage arises from the ability of MIGE to flexibly combine multiple reference entities in multimodal instructions and integrate additional visual features through its feature fusion mechanism.
-Methods DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
-Task-specific Models
-KOSMOS-G Pan et al. (2023) 0.694 0.847 0.287
-UNIMO-G Li et al. (2024b) 0.668 0.841 0.329
-Universal Models
-Omnigen (our test) Xiao et al. (2024b) 0.711 0.800 0.312
-Unireal Chen et al. (2024) 0.702 0.806 0.326
-MIGE (ours) 0.744 0.830 0.293
-          /only_subject data 0.726 0.823 0.289
-          /wo_VAE feature 0.741 0.828 0.287
-Table 1: Quantitative results for subject-driven image generation on DreamBench. MIGE outperforms universal models in subject preservation and remains competitive with task-specific models.
-4.2.2 Instruction-based Image Editing
-Instruction-based image editing enables users to modify a source image based on free-form multimodal instructions, including adding, removing, modifying object properties, or altering the overall style.
-Table 2 presents a quantitative analysis of the Emu Edit and MagicBrush test sets.
-DINO and CLIP-I evaluate similarity to the source image, while CLIP-T measures alignment with the target description.
-CLIPdir quantifies the alignment between CLIP text and image embedding changes, whereas L1 and L2 capture pixel-level differences.
-As shown in Table 2, MIGE achieves the highest CLIP-T score and surpasses all task-specific models in CLIPdir, demonstrating its superior ability to effectively follow multimodal instructions.
-As a universal model, MIGE outperforms all other universal models across all metrics on the MagicBrush test set, achieving the lowest L1 and L2 scores and the highest CLIP-I, DINO, and CLIP-T scores, highlighting its strong instruction fidelity and fine-detail preservation.
-This capability is further illustrated in Figure 5, where MIGE is the only model that accurately follows the instruction to add a Daffy Duck image to the red suitcase without altering other unrelated areas.
+During training, we sample instruction-based subject addition and subject replacement data in a 1:1 ratio.
+Separately, subject-driven editing and instruction-based editing data are also sampled in a 1:1 ratio.
+A training strategy is employed with a 5% probability of dropping either the conditional input or the multimodal condition, and an additional 5% chance of dropping both, enabling classifier-free guidance during inference.
+4.2. Evaluation Results
+As a unified model, MIGE demonstrates exceptional performance across various image generation and editing tasks, even outperforming existing task-specific models.
+MIGE achieves strong performance in subject-driven generation and instruction-based editing, and shows promising results on instruction-based subject-driven image generation.
+Detailed results are in Supplementary Material.
+Table 1. Quantitative results on the Emu Edit and MagicBrush test sets. MIGE outperforms other universal models on MagicBrush, demonstrating superior instruction alignment and detail preservation.
 Emu Edit Test set MagicBrush Test set
 […]
-Task-specific Models
-InstructPix2Pix Brooks et al. (2023) 0.759 0.831 0.288 0.086 0.122 0.036 0.763 0.843 0.289 0.105 0.097 0.028
-MagicBrush Zhang et al. (2024) 0.800 0.857 0.295 0.102 0.085 0.027 0.847 0.888 0.304 0.127 0.062 0.018
-UltraEdit Zhao et al. (2024) 0.862 0.875 0.301 0.100 0.050 0.008 0.879 0.897 0.305 0.119 0.042 0.006
-Universal Models
-PixWizard Lin et al. (2024) 0.824 0.862 0.288 0.066 0.105 0.038 0.865 0.897 0.289 0.054 0.073 0.023
-ACE Han et al. (2024) 0.847 0.877 0.299 0.104 0.079 0.028 0.856 0.899 0.304 0.127 0.065 0.023
-OmniGen Xiao et al. (2024b) 0.766 0.819 0.299 0.123 0.160 0.065 0.821 0.863 0.289 0.087 0.116 0.045
+Task-specific Models InstructPix2Pix (Brooks et al., 2023) 0.759 0.831 0.288 0.086 0.122 0.036 0.763 0.843 0.289 0.105 0.097 0.028
+MagicBrush (Zhang et al., 2024) 0.800 0.857 0.295 0.102 0.085 0.027 0.847 0.888 0.304 0.127 0.062 0.018
+UltraEdit (Zhao et al., 2024) 0.862 0.875 0.301 0.100 0.050 0.008 0.879 0.897 0.305 0.119 0.042 0.006
+SmartEdit (Huang et al., 2024b) 0.752 0.830 0.290 0.100 0.129 0.055 0.821 0.879 0.301 0.122 0.084 0.033
+ml-MGIE (Fu et al., 2023) 0.779 0.843 0.276 0.040 0.101 0.034 0.751 0.844 0.274 0.039 0.102 0.034
+Unified Models PixWizard (Lin et al., 2024) 0.824 0.862 0.288 0.066 0.105 0.038 0.865 0.897 0.289 0.054 0.073 0.023
+ACE (Han et al., 2024) 0.847 0.877 0.299 0.104 0.079 0.028 0.856 0.899 0.304 0.127 0.065 0.023
+OmniGen (Xiao et al., 2024) 0.766 0.819 0.299 0.123 0.160 0.065 0.821 0.863 0.289 0.087 0.116 0.045
+OneDiff (Le et al., 2024) 0.728 0.819 0.285 0.082 0.136 0.060 0.742 0.840 0.295 0.107 0.128 0.052
 MIGE (ours) 0.832 0.865 0.306 0.114 0.088 0.027 0.889 0.905 0.306 0.119 0.055 0.013
@@
 /wo_multimodal instruction 0.592 0.723 0.280 0.113 0.177 0.074 0.548 0.714 0.277 0.135 0.170 0.066
-Table 2: Quantitative results for instruction-based image editing on Emu Edit test set and MagicBrush test set.
-MIGE achieves the best overall performance on MagicBrush test set among universal models, with strong instruction fidelity and detail preservation.
-Figure 6: Qualitative results on the benchmark for the subject addition and subject replacement.
-The upper section compares subject addition results, while the lower section compares subject replacement.
-During testing, the <imagehere> placeholder in the multimodal instruction is replaced according to the image sequence.
-MIGE demonstrates flexibility in editing and excels in subject preservation ability and input-output consistency.
-4.2.3 Instruction-based Subject-driven
-   Image Editing
-Benchmark Construction
-Instruction-based subject-driven image editing is a novel task.
-Existing methods rely on masks or positional coordinates for editing Li et al.
-(2023b); Yang et al.
-(2024) but lack support for instruction-based editing.
-Current benchmarks for subject addition and replacement separately evaluate foreground and background similarities without providing a complete edited image as ground truth, making them unsuitable for this task.
-To address these issues, we construct a benchmark of 1,000 test samples: 500 for instruction-based subject addition and 500 for subject replacement.
-Data is sourced from SEED-Data-Edit Ge et al.
-(2024), with subjects extracted using Grounded SAM Ren et al.
-(2024).
-Captions for target images are generated by GPT-4o and refined through manual review.
-For compatibility with prior methods, our benchmark also includes masks.
-Details of the construction process are provided in Appendix A.
-Evaluation Results
-Our evaluation focuses on editing ability and subject preservation.
-Editing ability is assessed using DINO, CLIP-I, and CLIP-T: DINO and CLIP-I measure similarity to the ground truth image, while CLIP-T evaluates alignment with the target caption.
-Subject preservation is evaluated by extracting the edited subject via Grounded SAM and comparing it to the input subject image using DINO and CLIP-I.
-This separates the evaluation of image-level editing from subject-level feature preservation.
-Methods that do not support instruction-based editing are marked with a cross in the table, and masks used during testing.
-Quantitative comparisons with other methods are shown in Tables 3(a) and 3(b).
-We computed DINO and CLIP-I metrics between the source and target images, as well as the CLIP-T metric between the source image and target caption, shown in the first row of the table (denoted as “source-target”) as a baseline.
-In terms of editing ability, MIGE achieves the highest overall improvement across all metrics, demonstrating its effectiveness in performing edits guided by multimodal instructions.
-As shown in the qualitative comparison in Figure 6, MIGE understands the meaning of “replace” in instructions, rather than simply pasting the entity onto the image.
-For subject preservation, the results show that MIGE achieves the best performance in both tasks, as demonstrated in Figure 6.
+4.2.1. Instruction-based Image Editing
+Instruction-based image editing allows users to modify a source image using free-form multimodal instructions, such as adding, removing, or altering objects and styles.
+Table 1 reports results on Emu Edit and MagicBrush (Zhang et al., 2024).
+DINO and CLIP-I assess similarity to the source image, CLIP-T evaluates alignment with the target caption, CLIPdir measures consistency between text and image embedding changes, and L1/L2 capture pixel-level differences.
+As shown in Table 1, MIGE achieves the highest CLIP-T and outperforms all task-specific models on CLIPdir, indicating stronger instruction following.
+On MagicBrush, it surpasses all universal models across almost all metrics, with the best DINO, CLIP-I, CLIP-T, and lowest L1/L2, demonstrating superior instruction fidelity and detail preservation.
+This capability is further illustrated in Figure 6, where MIGE is the only model that accurately follows the instruction to add a Daffy Duck image to the red suitcase without altering other unrelated areas.
+Compared to ml-MGIE, which also uses a MLLM as an encoder, MIGE better preserves input-output consistency and avoids altering irrelevant backgrounds.
+Figure 6. Qualitative comparison for instruction-based image editing. MIGE demonstrates superior editing accuracy and instruction understanding.
+4.2.2. Subject-driven Image Generation
+Generating images that satisfy both image and text constraints from a multimodal prompt is challenging.
+We compare MIGE with three task-specific methods using MLLMs as encoders and three universal models on DreamBench (Ruiz et al., 2023).
+DINO and CLIP-I measure subject fidelity, while CLIP-T assesses instruction adherence.
+As shown in Table 2, MIGE surpasses all unified models in subject preservation ability, as evidenced by the highest DINO and CLIP-I scores, with the DINO score even outperforming task-specific models.
+Examples in Figure 7 further highlight MIGE’s ability to effectively preserve the identity of each object in multi-object scenarios and combine them based on instructions, demonstrating that our multimodal encoder effectively captures both semantic and visual features simultaneously.
+Table 2.
+Quantitative results for subject-driven image generation on DreamBench.
+Results with * denote our reproduction.
+MIGE outperforms universal models in subject preservation and is competitive with task-specific models.
+Methods DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
+Task-specific Models BLIP-Diffusion (Li et al., 2024a) 0.670 0.805 0.302
+KOSMOS-G (Pan et al., 2023) 0.694 0.847 0.287
+UNIMO-G (Li et al., 2024c) 0.668 0.841 0.329
+Unified Models Omnigen * (Xiao et al., 2024) 0.711 0.800 0.312
+UniReal (Chen et al., 2024c) 0.702 0.806 0.326
+OneDiffusion * (Le et al., 2024) 0.582 0.750 0.240
+MIGE (ours) 0.744 0.830 0.293
+/only_subject data 0.726 0.823 0.289
+/wo_VAE feature 0.741 0.828 0.287
+Figure 7. Qualitative comparison for subject-driven image generation. MIGE exhibits superior subject preservation, even in multi-subject scenarios.
+Table 3.
+Quantitative results on instruction-based subject-driven editing.
+Methods marked with a cross in the instruction column use masks, while the others generate images using multimodal instructions.
+Overall, MIGE significantly outperforms others in both tasks, showcasing superior editing and subject preservation abilities.
+(a) Results on instruction-based subject replacement.
 Methods Editing Subject Preserving
-Instruction DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
-source-target 0.668 0.842 0.271
-PBE Yang et al. (2023) 0.810 0.885 0.304 0.521 0.792
-TIGIC Li et al. (2025) 0.789 0.874 0.313 0.453 0.744
-MADD He et al. (2024) 0.736 0.852 0.284 0.446 0.742
-OmniGen Xiao et al. (2024b) 0.802 0.868 0.296 0.580 0.792
+DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
+I-O sim 0.668 0.842 0.271
+Mask-denpendent Models PBE (Yang et al., 2023) 0.810 0.885 0.304 0.521 0.792
+AnyDoor (Chen et al., 2024b) 0.817 0.890 0.307 0.551 0.789
+TIGIC (Li et al., 2025c) 0.789 0.874 0.313 0.453 0.744
+MADD (He et al., 2024a) 0.736 0.852 0.284 0.446 0.742
+Mask-free Models OmniGen (Xiao et al., 2024) 0.802 0.868 0.296 0.580 0.792
 MIGE(ours) 0.863 0.909 0.307 0.652 0.834
@@
 /only_compositional data 0.780 0.860 0.303 0.585 0.808
-((a)) Results on instruction-based subject replacement.
+/wo_VAE feature 0.852 0.900 0.308 0.633 0.819
+(b) Results on instruction-based subject addition.
 Methods Editing Subject Preserving
-Instruction DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
-source-target 0.783 0.903 0.295
-PBE Yang et al. (2023) 0.843 0.908 0.321 0.495 0.794
-TIGIC Li et al. (2025) 0.840 0.901 0.325 0.455 0.753
-MADD He et al. (2024) 0.885 0.930 0.316 0.519 0.785
-OmniGen Xiao et al. (2024b) 0.791 0.870 0.312 0.605 0.814
+DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$ DINO $\uparrow$ CLIP-I $\uparrow$
+I-O sim 0.783 0.903 0.295
+Mask-denpendent Models PBE (Yang et al., 2023) 0.843 0.908 0.321 0.495 0.794
+AnyDoor (Chen et al., 2024b) 0.863 0.930 0.324 0.533 0.799
+TIGIC (Li et al., 2025c) 0.840 0.901 0.325 0.455 0.753
+MADD (He et al., 2024a) 0.885 0.930 0.316 0.519 0.785
+Mask-free Models OmniGen (Xiao et al., 2024) 0.791 0.870 0.312 0.605 0.814
 MIGE(ours) 0.909 0.940 0.322 0.638 0.838
@@
 /only_compositional data 0.879 0.928 0.324 0.577 0.820
-((b)) Results on instruction-based subject addition.
-Table 3: Quantitative results on instruction-based subject-driven editing.
-Methods marked with a cross in the instruction column use masks, while the others generate images using multimodal instructions.
-Overall, MIGE significantly outperforms others in both tasks, showcasing superior editing and subject preservation abilities.
-4.3 Further Analysis
-4.3.1 Effectiveness of Joint Training
+/wo_VAE feature 0.902 0.934 0.322 0.633 0.831
+4.2.3. Instruction-based Subject-driven Image Editing
+Benchmark Construction
+Instruction-based subject-driven image editing is a novel task.
+Existing methods rely on masks or positional cues for editing (Li et al., 2023a; Yang et al., 2024) but lack support for instruction-only editing.
+The AnyEdit-Test Benchmark (Yu et al., 2024) supports few instruction templates and evaluates only one type of object editing based on automatically extracted masks.
+Current benchmarks (Yang et al., 2023; Chen et al., 2024b) for subject addition and replacement separately evaluate foreground and background similarities without providing a complete edited image as ground truth, making them unsuitable for this task.
+To address these issues, we manual construct MIGEBench, a benchmark consisting of 1,000 test samples: 500 for instruction-based subject addition and 500 for subject replacement.
+More details can be found in the Supplementary Material.
+Evaluation Results
+Our evaluation focuses on editing ability and subject preservation.
+DINO and CLIP-I measure the similarity of the generated image to the ground truth, while CLIP-T evaluates alignment with the target caption, indicating editing ability.
+Subject preservation is evaluated by extracting the edited subject via Grounded SAM (Ren et al., 2024) and comparing it to the input subject image using DINO and CLIP-I.
+This separates the evaluation of image-level editing from subject-level feature preservation.
+Methods that do not support instruction-based editing use masks provided in MIGEBench during testing.
+Quantitative comparisons with other methods are shown in Tables 3(a) and 3(b).
+We compute the DINO, CLIP-I, and CLIP-T scores between the input image and the ground truth.
+The first row of the table (denoted as ”I-O sim”) provides a baseline for evaluating the effectiveness of editing ability.
+Mask-dependent models can easily preserve the edit-unrelated regions, but MIGE still outperforms all methods, highlighting the effectiveness of conditional inputs and MIGE’s precise understanding of instructions.
+The CLIP-T score measures instruction adherence in mask-free models.
+MIGE significantly outperforms OmniGen in both tasks, demonstrating the advantage of our unified approach.
+MIGE also achieves the best performance on subject preservation in both tasks.
+As shown in Figure 8, MIGE successfully understands the task of replacing a bird with an owl, rather than simply overlaying or adding an owl, demonstrating the effectiveness of the multimodal encoder which simultaneously encoding the semantic and visual information of the reference image.
+Figure 8.
+Qualitative results on the benchmark for the subject addition and subject replacement.
+The upper section compares subject addition results, while the lower section compares subject replacement.
+During testing, the ¡imagehere¿ placeholder in the multimodal instruction is replaced according to the image sequence.
+MIGE demonstrates flexibility in editing and excels in subject preservation ability and input-output consistency.
+4.3. Further Analysis
+4.3.1. Effectiveness of Joint Training
 To assess the effectiveness of joint training, we train models separately on individual datasets (denoted as “only_subject data,” “only_edit data,” and “only_compositional data”) and compare their performance with the jointly trained model.
-The results in Table 1 and Table 2 show that joint training leads to consistent improvements across all metrics, demonstrating that subject-driven generation and instruction-based editing reinforce each other.
+The results in Tables 1, 2 and 3 show that joint training leads to consistent improvements across all metrics, demonstrating that subject-driven generation and instruction-based editing reinforce each other.
 As shown in Table 3, joint training also enhances the performance of the compositional new task, further highlighting its overall benefits.
@@
 In conclusion, joint training of subject-driven generation and instruction-based editing within our unified framework not only boosts compositional capability but also improves the performance of each individual task.
-4.3.2 Effectiveness of Feature Fusing
+This validates our motivation and provides further theoretical support for unified models for multiple tasks.
+4.3.2. Effectiveness of Feature Fusing
 MIGE employs a feature fusion mechanism in the multimodal encoder to integrate semantic features from ViT and visual features from VAE.
-As shown in Table 1 and Table 2, compared to the model without VAE features (denoted as “wo_VAE feature”), incorporating VAE features significantly improves detail preservation in reference images, benefiting both subject-driven image generation and instruction-based image editing.
+As shown in Table 2, Table 1 and Table 3, compared to the model without VAE features (denoted as “wo_VAE feature”), incorporating VAE features significantly improves detail preservation in reference images, benefiting both subject-driven image generation and instruction-based image editing.
 This is particularly evident in the improved CLIP-I and DINO scores and the significant reduction in L1 and L2 metrics, demonstrating that the inclusion of additional visual features helps maintain consistency between the input and output.
-4.3.3 Effectiveness of Instruction-based Subject-driven Image Editing Dataset
-Joint training in subject-driven image generation and instruction-based image editing enables generalization to instruction-based subject-driven image editing (denoted as “subject data + edit data”).
+4.3.3. Effectiveness of Instruction-based Subject-driven Image Editing Dataset
+Joint training of subject-driven image generation and instruction-based image editing enables generalization to instruction-based subject-driven image editing (denoted as “subject data + edit data”).
 To enhance the capability of MIGE in this new task, particularly in understanding spatial terms and size descriptions, we constructed a task-specific dataset for joint training.
@@
 This demonstrates the effectiveness of our constructed dataset, and the proposed data generation pipeline serves as a valuable reference for future dataset construction.
-4.3.4 Effectiveness of Multimodal Instruction
-Existing instruction-based editing works Brooks et al.
-(2023); Sheynin et al.
-(2024); Zhang et al.
-(2024) typically use text instructions as the conditional input, while we extend this to multimodal instructions.
+4.3.4. Effectiveness of Multimodal Instruction
+Existing instruction-based editing works (Brooks et al., 2023; Sheynin et al., 2024; Zhang et al., 2024) typically use text instructions as the conditional input, while we extend this to multimodal instructions.
 To measure the benefit of multimodal instructions, we trained the model with text-only editing instructions for comparison.
+As a unified model, MIGE demonstrates exceptional performance across various image generation and editing tasks, even outperforming existing task-specific models.
 As shown in Table 1, using multimodal instructions consistently improves performance over text-only instructions (denoted as “wo_multimodal instruction”).
@@
 While text-only instructions provide the necessary changes, the high CLIPdir score and lower values in other metrics show that multimodal instructions add visual context, enabling more precise and faithful modifications.
-5 Conclusion
-We present MIGE, a unified framework that combines subject-driven generation and instruction-based editing.
-By leveraging multimodal instructions and conditional input, MIGE enables joint training that reinforces both tasks.
-This approach improves task synergy and addresses data scarcity.
+For more qualitative results, see Supplementary Material.
+4.3.5. Comparison of Data Volume
+Task-specific models treat tasks separately and typically require large amounts of high-quality data.
+In contrast, by exploiting the relationships between tasks and the unified structure, MIGE achieves excellent results with minimal data through joint training.
+As shown in Table 4, MIGE demonstrates that both the model structure and joint training are key to achieving this performance with minimal data, using fewer data than task-specific models and other unified models.
+Table 4. Comparison of Data Volume Across Methods
+      Methods       Data Volume
+             Task-specific       Models       MADD (He et al., 2024a)       3M
+      PBE (Yang et al., 2023)       16M
+      KOSMOS-G (Pan et al., 2023)       3159.3M
+      UNIMO-G (Li et al., 2024c)       304.3M
+             Unified       Models       PixWizard (Lin et al., 2024)       30M
+      UniReal (Chen et al., 2024c)       100M
+      ACE (Han et al., 2024)       700M
+      Omnigen (Xiao et al., 2024)       326.06M
+      MIGE (ours)       2.28M
+5. Conclusion
+We present MIGE, a unified framework that combines subject-driven generation and instruction-based editing, leveraging multimodal instructions and conditional input for joint training that enhances task synergy.
 Joint training also unlocks new capabilities like instruction-based subject-driven image editing.
 We introduce pipelines for this new task to construct training data and MIGEBench for evaluation.
-Our experiments show that joint training leads to significant improvements in subject fidelity and instruction adherence, demonstrating the effectiveness of unifying these tasks.
+Our experiments show that joint training leads to significant enhancement in subject fidelity and instruction adherence, demonstrating the effectiveness of unifying these tasks.
 This integration enhances controllability and offers promising directions for future multimodal image generation and editing.
-6 Limitations
-Our model faces common challenges in context accuracy, complex composition, and visual faithfulness, especially in multi-entity tasks.
-Ethical concerns, like deepfakes, also arise.
-Nevertheless, our framework shows promise in enabling unified controlled image generation.
-Instruction-based subject-driven generation remains challenging, with difficulties in handling spatial relationships and adjusting subject sizes to fit the background.
-Additionally, the technology carries the risk of generating misleading images, but our work is intended solely for research, emphasizing its exploratory nature rather than real-world deployment.
-Appendix A Benchmark Construction
-Instruction-based subject-driven image editing is a novel task lacking evaluation benchmarks.
-Previous works Li et al.
-(2023b); Yang et al.
-(2024) did not support instructions and lacked ground truth target images.
-To address this, we designed a benchmark construction pipeline with multiple manual inspections to ensure image and caption quality.
-As shown in Figure 7, we first filter valid editing pairs using a Gradio Abid et al.
-(2019) interface, as shown in Figure 8.
-Then, Qwen2.5-14B-Instruct extracts entity names and constructs multimodal instructions.
-After manually reviewing object names, we use Grounding DINO Liu et al.
-(2025) for bounding boxes and SAM for segmentation.
-Cropped entities are saved with black and white backgrounds for diverse test scenarios.
-Prompts are detailed in Appendix E.
-For evaluation, Qwen2-VL-7B generates captions that compare the generated images with their ground truth counterparts.
-These captions are then manually reviewed to identify and remove any unrecognized entities.
-This review process ensures that the generated results are accurate and reliable.
-The final benchmark consists of 500 samples for both subject addition and replacement tasks.
-By carefully avoiding artifacts from inpainting and incorporating multiple rounds of manual review, we guarantee the high quality of the data.
-Figures 9 and 10 display some example pairs from the benchmark to illustrate the evaluation process.
-Figure 7: The pipeline of benchmark construction.
-Figure 8: The annotation interface for benchmarks is built with Gradio.
-Figure 9: Subject addition examples in our benchmark.
-Figure 10: Subject replacement examples in our benchmark.
-Appendix B Dataset Construction
-To enable joint training of MIGE, we designed a series of pipelines for data construction and processing.
-Our dataset includes three tasks: subject-driven image generation, instruction-based image editing, and instruction-based subject-driven image editing.
-The proportions of the training data are shown in Figure 11.
-In this section, we will detail the pipelines for processing each data component and other related details.
-Figure 11: Composition of training data.
-Subject-driven Image Generation Data Construction
-We construct multimodal instruction-image pairs by replacing entities in image captions with corresponding images.
-The dataset is sourced from BLIP3-GROUNDING-50M Xue et al.
-(2024) and internal data.
-For the former, we discard low-quality pairs with CLIP-T scores below 0.255.
-Entities are filtered based on annotated bounding boxes, retaining those with sizes between 0.05 and 0.8.
-After segmenting entities using SAM Kirillov et al.
-(2023), we further compute the CLIP-T score between “a photo of [entity]” and the segmented image to mitigate incomplete pairs.
-For internal data, we follow the KOSMOS-G and UNIMO-G processing pipelines, using MPT-7B-Instruct Team et al.
-(2023) to extract entities.
-The original image and entity text are input into Grounded SAM to obtain entity images, which are filtered by a CLIP-T score above 0.255 and a size ratio between 0.1 and 0.85.
-To enhance subject preservation, we incorporate 112,846 samples from subject200k, as in OminiControl Tan et al.
-(2024).
-Instruction-based Image Editing Data Construction
-Previous works on instruction-based editing relied solely on text, covering object-level additions, deletions, modifications, and global changes in background or style.
-We integrated data from InstructPix2Pix Brooks et al.
-(2023), UltraEdit Zhao et al.
-(2024), MagicBrush Zhang et al.
-(2024), SEED-Data-Edit Ge et al.
-(2024), GIER Shi et al.
-(2020), and HQ-Edit Hui et al.
-(2024), filtering out instances with target image aesthetic scores above 5.5.
-To enhance multimodal instruction alignment, we reformulated text-based instructions based on task type.
-Specifically, for instructions containing phrases like “in the image” or “to the image,” we replaced them with “in<imagehere>” or “to<imagehere>.” For others, we appended “of<imagehere>” for global edits or “in<imagehere>” for localized modifications.
-Instruction-based Subject-driven Image Editing Data Construction
-To enhance instruction-based subject-driven image editing, we propose dedicated data construction pipelines.
-For subject addition, we design a pipeline inspired by SAM-FB He et al.
-(2024) to construct background images, foreground entities, and multimodal instructions, as shown in Figure 4(a).
-We start with filtering SA-1B Kirillov et al.
-(2023) images with an aesthetic score above 5 and select entities where the foreground-to-image ratio is between 0.1 and 0.5.
-Qwen2-VL-7B Wang et al.
-(2024) is used to verify entity completeness.
-The selected entities serve as foregrounds, while the backgrounds are inpainted using LAMA Suvorov et al.
-(2022).
-We then extract foreground-related text using Qwen2-VL-7B and generate multimodal instructions with GPT-4o Hurst et al.
-(2024).
-Due to resource constraints, we process the first 500,000 SA-1B samples, yielding 193,247 multimodal instruction-foreground-background pairs.
-For subject replacement, as shown in Figure 4(b), we aim at constructing foreground entities and multimodal instructions.
-We filter SEED-Data-Edit’s part1-unsplash based on predefined rules, retaining only replace-task data where the match score in the annotations exceeds 0.3 and the target image’s aesthetic score is above 5.5.
-Qwen2.5-14B-Instruct Team (2024) identifies whether the main entity is being replaced.
-The target entity is replaced with <imagehere> to construct a multimodal prompt.
-Using the annotated edited subject information, we apply Grounded SAM for segmentation, obtaining 79,693 multimodal instruction-replace entity image-source image-target image pairs.
-In addition, we construct 34,947 virtual try-on samples using IDM-VTON Choi et al.
-(2024).
-Appendix C Evaluation Details
-We evaluated the subject-driven generation capability of MIGE on DreamBench, which contains 750 prompts covering 30 subjects.
-For each prompt, four images were generated using seeds 0, 1, 2, and 3, resulting in a total of 3,000 images.
-Following KOSMOS-G, we selected one image per prompt.
-Then we extracted the subject using Grounded SAM and used it as input, aligning with the training process.
-Subject fidelity was assessed using DINO and CLIP-I, while CLIP-T measured adherence to multimodal instructions.
-For instruction-based editing evaluation, the Emu Edit benchmark contains known issues, including incorrect image-caption pairs and duplicate source-target captions.
-Prior works have handled these inconsistencies differently, leading to incomparable results.
-Therefore, we reimplemented all currently available open-source methods on both test sets using a fixed random seed (seed=0) for consistency.
-Consequently, CLIP-I, L1, and DINO were computed on 3,589 Emu Edit test samples, while CLIPdir and CLIP-T were evaluated on 2,899 samples after removing problematic data.
-Appendix D Related Work
-Subject-Driven Image Generation
-Recent advances in generative models have significantly improved image customization, enabling finer and more flexible control over generated content.
-Subject-driven image generation, which preserves the features of given images while creating new content, has garnered increasing attention.
-DreamBooth Ruiz et al.
-(2023) fine-tunes the entire U-Net using multiple images of the same subject, making it resource-intensive.
-IP-Adapter Ye et al.
-(2023) introduces a test-time tuning-free approach by decoupling image and text, training only additional attention layers, but struggles to balance prompt adherence and image fidelity.
-FastComposer Xiao et al.
-(2024a) and Blip-Diffusion Li et al.
-(2024a) leverage embeddings to condition images on text but face challenges in subject preservation and efficiency with generalized vision-language inputs.
-KOSMOS-G Pan et al.
-(2023) and UNIMO-G Li et al.
-(2024b) use Multimodal Large Language Models (MLLMs) as multimodal encoders for greater input flexibility but require large datasets to train task-specific encoders and alignment networks.
-However, these methods lack strong instruction-following capabilities, limiting their scalability to new tasks and applications.
-Instruction-based Image Editing
-Instruction-based image editing modifies images at the object level (e.g., adding, removing, or altering objects) or the image level (e.g., changing style or environment) based on user instructions.
-This approach offers greater flexibility than caption-based methods but demands stronger instruction-following capabilities.
-Inst-inpaint Yildirim et al.
-(2023) focuses on object removal, while EraseDraw Canberk et al.
-(2025) fine-tunes a model for object addition using data derived from removal tasks.
-InstructPix2Pix Brooks et al.
-(2023), trained on a large synthetic dataset, supports four types of edits, and MagicBrush Zhang et al.
-(2024) refines this capability by fine-tuning on real, manually annotated data.
-EmuEdit Sheynin et al.
-(2024) enhances diffusion-based editing with a task vector to differentiate tasks.
-However, these methods often struggle with maintaining input-output consistency.
-Figure 12: Qualitative results of subject-driven image generation (top) , instruction-based image editing (middle), and instruction-based subject-driven image editing (bottom).
-Appendix E Prompts
-LLMs and MLLMs are integral to our training data and benchmark construction pipeline, and the prompts used are presented in this section.
-Prompt for Qwen2-VL-7B to determine the completeness of subject in the image
-Prompt: Determine if the subject in the image is complete.
-If it is complete and not an abstract object such as background, grass, sky, tree, stone, or part of another item, please return True.
-Otherwise, return False.
-Table 4: Prompt design for subject completeness evaluation in an image. The model should identify whether the subject is complete, not abstract elements.
-Prompt for GPT-4o to generate multimodal instruction
-Prompt: The object is {object_name}.
-It is located in a bounding box with coordinates ({x}, {y}, {w}, {h}) on an image of size {width}x{height}.
-Describe its size, relative position, and relation to surrounding objects.
-Avoid describing the overall scene or unrelated elements.
-Your response should start with “Add <imagehere> to the [position] of <imagehere>”.
-(The first <imagehere> indicates the object and the second indicates the image.) Keep the <imagehere> symbol in the first sentence in your reply.
-Answer briefly in two sentences:
-Table 5: Prompt design for GPT-4o to generate multimodal instruction for subject addition training data.
-Prompt for Qwen2.5-14B-Instruct to determine whether an edit instruction pertains to the main subject of an image
-Prompt: You are an assistant that determines whether an edit instruction pertains to the main subject of an image.
-The main subject refers to humans or animals only. If the edit instruction is related to the main subject, respond with ’yes’.
-If it pertains to background, large areas of vegetation, or environmental information, respond with ’no’.
-Here are some examples:
-1. replace the grass with sand $\rightarrow$ no
-2. replace the trees with palm trees $\rightarrow$ no
-3. replace the dirt road with a cobblestone path $\rightarrow$ no
-4. replace the bird with a parrot $\rightarrow$ yes
-Now, analyze the following instruction and respond accordingly: {instruction}
-Table 6: Prompt design for determining if an edit instruction pertains to the main subject of the image.
-The model evaluates whether the instruction relates to humans, animals, or other environmental features.
-Prompt for extracting the main subject and modifying the edit instruction
-Prompt 1: Replace Examples
-You are an assistant that determines the main subject of an edit instruction and outputs the extracted main subject along with a modified prompt.
-Given an edit instruction like “replace X with Y”:
-- Extract “Y” (the replacement subject).
-- Output the extracted subject.
-- Modify the instruction by replacing “Y” with <imagehere> in the original prompt and output the new instruction.
-Examples:
-1. replace one woman with a man $\rightarrow$ Output: a man, replace one woman with <imagehere>
-2. replace the castle with another castle $\rightarrow$ Output: castle, replace the castle with <imagehere>
-3. replace the desk with a white one in the bottom middle $\rightarrow$ Output: desk, replace the desk with <imagehere> in the bottom middle
-Prompt 2: Add Examples
-You are an assistant that extracts the added subject in the edit instruction.
-Given an edit instruction like “add X to the image”:
-- Extract “X” (the added subject).
-- Output the extracted subject.
-- Modify the instruction by replacing “X” with <imagehere> in the original prompt and output the new instruction.
-Here are some examples:
-1. add human over the stone in the bottom left $\rightarrow$ Output: human, add <imagehere> over the stone in the bottom left
-2. add a car on the road at the bottom $\rightarrow$ Output: a car, add <imagehere> on the road at the bottom
-3. Add an owl on the left shoulder $\rightarrow$ Output: an owl, Add <imagehere> on the left shoulder
-Table 7: Prompt designs for extracting and modifying edit instructions based on either replacement or addition of subjects.

### v4

@@
 Controllable Image Generation, Diffusion Model, Joint Training
-copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the 33th ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06submissionid: 6167ccs: Computing methodologies Image representations
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755811isbn: 979-8-4007-2035-2/2025/10submissionid: 6167ccs: Computing methodologies Image representations
 1. Introduction
@@
 4.1. Implementation Details
-MIGE includes a conditional diffusion model and a multimodal encoder.
-Our design allows flexible selection of various diffusion models, and we initialize with PIXART-$\alpha$ (Chen et al., 2023) pretrained at a 512×512 resolution.
-The multimodal encoder consists of a pretrained Flan-T5-XXL (Chung et al., 2024) as LLM for initialization and an image encoding component.
-This includes query tokens, Q-Former, and a projector, all initialized with a BLIP-2 (Li et al., 2023b) checkpoint (pretrain_flant5xxl).
-The frozen VAE encoder, used as the visual feature extractor, is the same as the one in the diffusion model.
-During training, we sample instruction-based subject addition and subject replacement data in a 1:1 ratio.
-Separately, subject-driven editing and instruction-based editing data are also sampled in a 1:1 ratio.
-A training strategy is employed with a 5% probability of dropping either the conditional input or the multimodal condition, and an additional 5% chance of dropping both, enabling classifier-free guidance during inference.
+[…]
 4.2. Evaluation Results
@@
 […]
-Task-specific Models InstructPix2Pix (Brooks et al., 2023) 0.759 0.831 0.288 0.086 0.122 0.036 0.763 0.843 0.289 0.105 0.097 0.028
+Task-specific
+Models
+InstructPix2Pix (Brooks et al., 2023) 0.759 0.831 0.288 0.086 0.122 0.036 0.763 0.843 0.289 0.105 0.097 0.028
 MagicBrush (Zhang et al., 2024) 0.800 0.857 0.295 0.102 0.085 0.027 0.847 0.888 0.304 0.127 0.062 0.018
@@
 ml-MGIE (Fu et al., 2023) 0.779 0.843 0.276 0.040 0.101 0.034 0.751 0.844 0.274 0.039 0.102 0.034
-Unified Models PixWizard (Lin et al., 2024) 0.824 0.862 0.288 0.066 0.105 0.038 0.865 0.897 0.289 0.054 0.073 0.023
+Unified
+Models
+PixWizard (Lin et al., 2024) 0.824 0.862 0.288 0.066 0.105 0.038 0.865 0.897 0.289 0.054 0.073 0.023
 ACE (Han et al., 2024) 0.847 0.877 0.299 0.104 0.079 0.028 0.856 0.899 0.304 0.127 0.065 0.023
@@
 Methods DINO $\uparrow$ CLIP-I $\uparrow$ CLIP-T $\uparrow$
-Task-specific Models BLIP-Diffusion (Li et al., 2024a) 0.670 0.805 0.302
+Task-specific
+Models
+BLIP-Diffusion (Li et al., 2024a) 0.670 0.805 0.302
 KOSMOS-G (Pan et al., 2023) 0.694 0.847 0.287
 UNIMO-G (Li et al., 2024c) 0.668 0.841 0.329
-Unified Models Omnigen * (Xiao et al., 2024) 0.711 0.800 0.312
+Unified
+Models
+Omnigen * (Xiao et al., 2024) 0.711 0.800 0.312
 UniReal (Chen et al., 2024c) 0.702 0.806 0.326
@@
 I-O sim 0.668 0.842 0.271
-Mask-denpendent Models PBE (Yang et al., 2023) 0.810 0.885 0.304 0.521 0.792
+Mask-denpendent
+Models
+PBE (Yang et al., 2023) 0.810 0.885 0.304 0.521 0.792
 AnyDoor (Chen et al., 2024b) 0.817 0.890 0.307 0.551 0.789
@@
 MADD (He et al., 2024a) 0.736 0.852 0.284 0.446 0.742
-Mask-free Models OmniGen (Xiao et al., 2024) 0.802 0.868 0.296 0.580 0.792
+Mask-free
+Models
+OmniGen (Xiao et al., 2024) 0.802 0.868 0.296 0.580 0.792
 MIGE(ours) 0.863 0.909 0.307 0.652 0.834
@@
 I-O sim 0.783 0.903 0.295
-Mask-denpendent Models PBE (Yang et al., 2023) 0.843 0.908 0.321 0.495 0.794
+Mask-denpendent
+Models
+PBE (Yang et al., 2023) 0.843 0.908 0.321 0.495 0.794
 AnyDoor (Chen et al., 2024b) 0.863 0.930 0.324 0.533 0.799
@@
 MADD (He et al., 2024a) 0.885 0.930 0.316 0.519 0.785
-Mask-free Models OmniGen (Xiao et al., 2024) 0.791 0.870 0.312 0.605 0.814
+Mask-free
+Models
+OmniGen (Xiao et al., 2024) 0.791 0.870 0.312 0.605 0.814
 MIGE(ours) 0.909 0.940 0.322 0.638 0.838
@@
   Methods   Data Volume
-             Task-specific       Models       MADD (He et al., 2024a)       3M
+  Task-specific
+  Models
+  MADD (He et al., 2024a)   3M
   PBE (Yang et al., 2023)   16M
@@
   UNIMO-G (Li et al., 2024c)   304.3M
-             Unified       Models       PixWizard (Lin et al., 2024)       30M
+  Unified
+  Models
+  PixWizard (Lin et al., 2024)   30M
   UniReal (Chen et al., 2024c)   100M
@@
 This integration enhances controllability and offers promising directions for future multimodal image generation and editing.
+6. Supplementary Material
+Appendix A Qualitative Comparison
+We conduct a qualitative comparison to assess the visual performance and effectiveness of MIGE.
+Figure 14 presents a range of cases on each task to showcase the generation results of MIGE, demonstrating its ability to handle diverse inputs and produce high-quality outputs.
+This comparison highlights the strengths of MIGE and the benefit of joint training, offering valuable insights into its potential for real-world applications.
+Appendix B Related Work
+B.1. Subject-Driven Image Generation
+Subject-driven generation focuses on preserving given subject features while generating new images.
+DreamBooth (Ruiz et al., 2023) fine-tunes the full model, leading to high resource costs.
+IP-Adapter (Ye et al., 2023) trains lightweight attention layers, struggles with prompt-image balance.
+Methods like Blip-Diffusion (Li et al., 2024a) use embeddings to condition generation, yet face challenges in subject preservation and generalization.
+KOSMOS-G (Pan et al., 2023) and UNIMO-G (Li et al., 2024c) leverage Multimodal Large Language Models (MLLMs) as flexible encoders, but require large datasets for task-specific adaptation and lack robust instruction-following ability, limiting transferability.
+B.2. Instruction-based Image Editing
+Instruction-based editing modifies images based on instructions, offering greater flexibility than caption-driven methods.
+Inst-inpaint (Yildirim et al., 2023) and EraseDraw (Canberk et al., 2025) focus on object-level edits.
+InstructPix2Pix (Brooks et al., 2023), EmuEdit (Sheynin et al., 2024), and MagicBrush (Zhang et al., 2024) enable diverse edits via large-scale data.
+AnyEdit (Yu et al., 2024), FireEdit (Zhou et al., 2025), and TF-TI2I (Hsiao et al., 2025) introduce task-aware routing, region prompts, or fixed templates.
+Recent works like InsightEdit (Xu et al., 2024), MGIE (Fu et al., 2023), and SmartEdit (Huang et al., 2024b) leverage MLLMs to enhance instruction understanding.
+However, maintaining input-output consistency remains challenging.
+Appendix C Dataset Construction
+To enable joint training of MIGE, we designed a series of pipelines for data construction and processing.
+Our dataset includes three tasks: subject-driven image generation, instruction-based image editing, and instruction-based subject-driven image editing.
+The proportions of the training data are shown in Figure 9.
+In this section, we will detail the pipelines for processing each data component and other related details.
+Figure 9. Composition of training data.
+C.1. Subject-driven Image Generation Data Construction
+We construct multimodal instruction-image pairs by replacing entities in image captions with corresponding entity images.
+The dataset combines BLIP3-GROUNDING-50M (Xue et al., 2024) and internal sources.
+For the former, we filter out low-quality pairs with CLIP-T scores below 0.255 and retain entities with bounding boxes sized between 0.05 and 0.8.
+Entities are segmented using SAM (Kirillov et al., 2023), and pairs with low CLIP-T scores between the prompt ”a photo of [entity]” and the segmented image are removed.
+For internal data, we adopt the KOSMOS-G’s pipeline, using MPT-7B-Instruct (Team et al., 2023) to extract entities.
+Entity images are obtained via Grounded SAM (Ren et al., 2024), and filtered by a CLIP-T score threshold of 0.255 and a size ratio between 0.1 and 0.85.
+To enhance subject fidelity, we further include 112,846 samples from subject200k, following OminiControl (Tan et al., 2024).
+C.2. Instruction-based Image Editing Data Construction
+Previous works on instruction-based editing relied solely on text, covering object-level additions, deletions, modifications, and global changes in background or style.
+We integrated data from InstructPix2Pix (Brooks et al., 2023), UltraEdit (Zhao et al., 2024), MagicBrush (Zhang et al., 2024), SEED-Data-Edit (Ge et al., 2024), GIER (Shi et al., 2020), and HQ-Edit (Hui et al., 2024), filtering out instances with target image aesthetic scores above 5.5.
+To enhance multimodal instruction alignment, we reformulated text-based instructions based on task type.
+Specifically, for instructions containing phrases like “in the image” or “to the image,” we replaced them with “in¡imagehere¿” or “to¡imagehere¿.” For others, we appended “of¡imagehere¿” for global edits or “in¡imagehere¿” for localized modifications.
+C.3. Instruction-based Subject-driven Image Editing Data Construction
+For subject addition, inspired by SAM-FB (He et al., 2024a), we construct background images, foreground entities, and multimodal instructions, as shown in Figure 5(a).
+We filter SA-1B (Kirillov et al., 2023) images with aesthetic scores above 5 and select entities with a foreground-to-image ratio between 0.1 and 0.5.
+Qwen2-VL-7B (Wang et al., 2024) verifies entity completeness.
+The selected entities serve as foregrounds, and backgrounds are inpainted using LAMA (Suvorov et al., 2022).
+Foreground-related text is extracted with Qwen2-VL-7B, and multimodal instructions are generated via GPT-4o (Hurst et al., 2024).
+Due to resource constraints, we process the first 500,000 SA-1B samples, yielding 193,247 multimodal instruction-foreground-background pairs.
+For subject replacement, we construct foreground entities and multimodal instructions, as shown in Figure 5(b).
+SEED-Data-Edit’s part1-unsplash is filtered to retain replace-task samples with annotation match scores above 0.3 and aesthetic scores above 5.5.
+Qwen2.5-14B-Instruct (Team, 2024) verifies the main entity.
+The target entity is replaced with <imagehere> to form the multimodal prompt.
+Grounded SAM segments entities based on annotations, resulting in 79,693 multimodal instruction-replace entity-source image-target image pairs.
+We also include 34,947 virtual try-on samples generated using IDM-VTON (Choi et al., 2024).
+Appendix D Prompts
+LLMs and MLLMs are essential to our data and benchmark construction, with all prompts listed in Table 6 to Table 9.
+Appendix E Implementation Details
+E.1. Training Details
+[…]
+MIGE is trained on our multi-task dataset using the AdamW optimizer (Loshchilov, 2017) with a weight decay of 0.03 and a learning rate of 1e-5 for 18 epochs on 48 H20 GPUs, totaling six days of training with a batch size of 960 (20 per GPU).
+E.2. Evaluation Details
+We evaluated the subject-driven generation capability of MIGE on DreamBench, which contains 750 prompts covering 30 subjects.
+For each prompt, four images were generated using seeds 0, 1, 2, and 3, resulting in 3,000 images.
+Following KOSMOS-G, we selected one image per prompt.
+Then we use the extracted subject from Grounded SAM as input, aligning with the training process.
+Subject fidelity was assessed using DINO and CLIP-I, while CLIP-T measured adherence to multimodal instructions.
+For instruction-based editing evaluation, the Emu Edit benchmark contains known issues, including incorrect image-caption pairs and duplicate source-target captions.
+Prior works have handled these inconsistencies differently, leading to incomparable results.
+Therefore, we reimplemented all currently available open-source methods on both test sets using a fixed random seed (seed=0) for consistency.
+Consequently, CLIP-I, L1, and DINO were computed on 3,589 Emu Edit test samples, while CLIPdir and CLIP-T were evaluated on 2,899 samples after removing problematic data.
+Figure 10. The pipeline of benchmark construction.
+Appendix F Benchmark Construction
+Instruction-based subject-driven image editing is a novel task lacking evaluation benchmarks.
+Previous works (Li et al., 2023a; Yang et al., 2024) did not support variety of instructions and lacked ground truth target images.
+To address this, we designed a benchmark construction pipeline with multiple manual inspections to ensure image and caption quality.
+As shown in Figure 10, we first filter valid editing pairs using a Gradio (Abid et al., 2019) interface, as shown in Figure 11.
+Then, Qwen2.5-14B-Instruct extracts entity names and constructs multimodal instructions.
+After manually reviewing object names, we use Grounding DINO (Liu et al., 2025) for bounding boxes and SAM for segmentation.
+Cropped entities are saved with black and white backgrounds for diverse test scenarios.
+Prompts are detailed in Supplementary Material D.
+For evaluation, Qwen2-VL-7B generates captions comparing generated and ground truth images, followed by manual reviews to remove unrecognized entities and correct inaccuracies.
+The final benchmark includes 500 samples each for subject addition and replacement.
+By avoiding artifacts from inpainting and incorporating multiple manual reviews, we ensure high-quality data.
+Figure 12 and Figure 13 showcase some example benchmark pairs.
+Figure 11. The annotation interface for benchmarks built with Gradio.
+Figure 12. Subject addition examples in our benchmark.
+Figure 13. Subject replacement examples in our benchmark.
+Our benchmark, MIGEBench, is constructed based on SEED-Data-Edit, which is diversity.
+The benchmark covers a total of 565 distinct subjects.
+For the subject-replacement subset, the data is distributed across 10 major, well-balanced categories, as shown in 5.
+This hierarchical and balanced structure ensures a high degree of diversity throughout the MIGEBench dataset.
+Table 5. Category-wise Statistics of MIGEBench
+Category Subject Replacement Subject Addition
+Animals 62 77
+People 60 66
+Vehicles 40 52
+Buildings & Structures 27 50
+Plants & Nature 30 33
+Household & Man-made Objects 88 91
+Art & Decor 54 47
+Food & Drink 23 8
+Clothing & Accessories 31 24
+Other 85 52
+Total 500 500
+Table 6. Prompt design for subject completeness evaluation in an image. The model should identify whether the subject is complete, not abstract elements.
+Prompt for Qwen2-VL-7B to determine the completeness of subject in the image
+Prompt: Determine if the subject in the image is complete.
+If it is complete and not an abstract object such as background, grass, sky, tree, stone, or part of another item, please return True.
+Otherwise, return False.
+Table 7. Prompt design for GPT-4o to generate multimodal instruction for subject addition training data.
+Prompt for GPT-4o to generate multimodal instruction
+Prompt: The object is {object_name}.
+It is located in a bounding box with coordinates ({x}, {y}, {w}, {h}) on an image of size {width}x{height}.
+Describe its size, relative position, and relation to surrounding objects.
+Avoid describing the overall scene or unrelated elements.
+Your response should start with “Add <imagehere> to the [position] of <imagehere>”.
+(The first <imagehere> indicates the object and the second indicates the image.) Keep the <imagehere> symbol in the first sentence in your reply.
+Answer briefly in two sentences:
+Table 8.
+Prompt design for determining if an edit instruction pertains to the main subject of the image.
+The model evaluates whether the instruction relates to humans, animals, or other environmental features.
+Prompt for Qwen2.5-14B-Instruct to determine whether an edit instruction pertains to the main subject of an image
+Prompt: You are an assistant that determines whether an edit instruction pertains to the main subject of an image.
+The main subject refers to humans or animals only. If the edit instruction is related to the main subject, respond with ’yes’.
+If it pertains to background, large areas of vegetation, or environmental information, respond with ’no’.
+Here are some examples:
+1. replace the grass with sand $\rightarrow$ no
+2. replace the trees with palm trees $\rightarrow$ no
+3. replace the dirt road with a cobblestone path $\rightarrow$ no
+4. replace the bird with a parrot $\rightarrow$ yes
+Now, analyze the following instruction and respond accordingly: {instruction}
+Table 9. Prompt designs for extracting and modifying edit instructions based on either replacement or addition of subjects.
+Prompt for extracting the main subject and modifying the edit instruction
+Prompt 1: Replace Examples
+You are an assistant that determines the main subject of an edit instruction and outputs the extracted main subject along with a modified prompt.
+Given an edit instruction like “replace X with Y”:
+- Extract “Y” (the replacement subject).
+- Output the extracted subject.
+- Modify the instruction by replacing “Y” with <imagehere> in the original prompt and output the new instruction.
+Examples:
+1. replace one woman with a man $\rightarrow$ Output: a man, replace one woman with <imagehere>
+2. replace the castle with another castle $\rightarrow$ Output: castle, replace the castle with <imagehere>
+3. replace the desk with a white one in the bottom middle $\rightarrow$ Output: desk, replace the desk with <imagehere> in the bottom middle
+Prompt 2: Add Examples
+You are an assistant that extracts the added subject in the edit instruction.
+Given an edit instruction like “add X to the image”:
+- Extract “X” (the added subject).
+- Output the extracted subject.
+- Modify the instruction by replacing “X” with <imagehere> in the original prompt and output the new instruction.
+Here are some examples:
+1. add human over the stone in the bottom left $\rightarrow$ Output: human, add <imagehere> over the stone in the bottom left
+2. add a car on the road at the bottom $\rightarrow$ Output: a car, add <imagehere> on the road at the bottom
+3. Add an owl on the left shoulder $\rightarrow$ Output: an owl, Add <imagehere> on the left shoulder
+Figure 14. Qualitative results of subject-driven image generation (top) , instruction-based image editing (middle), and instruction-based subject-driven image editing (bottom).
+Acknowledgements.
+This work was supported by the National Natural Science Foundation of China (No.62172393), and Major Public Welfare Project of Henan Province (No.201300311200).